### v1

# The Euclidean distance degree of curves: from rational to line multiview varieties

Abstract
The Euclidean distance (ED) degree is an invariant that measures the algebraic complexity of optimizing the distance function of a point to a model. It has been studied in algebraic statistics, machine learning, and computer vision. In this article, we prove a formula for the ED degree of curves parameterized by rational functions with mild genericity assumptions. We apply our results to resolve conjectures on one-dimensional line multiview varieties from computer vision proposed by Duff and Rydell.
Introduction
One of the most impactful applications of algebraic geometry is in computer vision. This has led to the development of what is now known as algebraic vision. A featured problem in this area is to determine the number of critical points of the triangulation and, more generally, the multiview triangulation problem in multiple view geometry [13]. This amounts to minimizing the Euclidean distance function to an algebraic set known as the (affine) multiview variety. Recent work by Duff and Rydell introduced an extensive catalogue of multiview varieties motivated by new vision applications. In turn, this led to questions and conjectures on the Euclidean distance degrees for these models. In this article, we prove two of these conjectures using a new result on the ED degree of curves in products of projective spaces.
Figure 1: A yellow cubic curve with its orange projection in the blue image plane (left). A yellow quartic curve with three orange projections in the blue image plane (right).
Our paper is structured as follows. In Section 1 we provide the necessary preliminaries for the tools we use to prove our main results. We also highlight prior work in the algebraic vision community to put our results into context and address some of the subtleties in the proofs. In Section 2 we prove a general statement on the Euclidean distance degree of the multiview variety of rational curves of arbitrary degree. In Section 3 we have the main application of our result: strengthening the catalogue theorem [7, Theorem 3.3] of Duff-Rydell by settling their conjectures on one-dimensional multiview varieties. In Section 4 we conclude with a brief discussion on future directions.
1 Preliminaries
1.1 Multiprojective varieties and multidegrees
A subvariety of a product of projective spaces is known as a multiprojective variety. Let $Z\subset(\mathbb{P}^{h})^{n}$ be such a variety. Its defining ideal lies in the multigraded ring
$\mathbb{K}[\overline{\mathbf{x}}_{1},\dots,\overline{\mathbf{x}}_{n}]\quad\text{ where }\quad\overline{\mathbf{x}}_{i}=(x_{i,0},\dots,x_{i,h})$
and $\mathbb{K}$ is a field. If $Z$ is irreducible and of dimension $\dim{Z}$, then the multidegree of $Z$ is the polynomial
$\operatorname{Deg}(Z):=\sum_{\alpha_{1}+\cdots+\alpha_{n}=\dim{Z}}c_{\alpha}\frac{T_{1}^{h}\cdots T_{n}^{h}}{T^{\alpha}}$
in the quotient ring $\mathbb{Z}[T_{1},\dots,T_{n}]/\langle T_{1}^{h+1},\dots,T_{n}^{h+1}\rangle$. We use the standard multi-index notation here, so that if $\alpha=(\alpha_{1},\dots,\alpha_{n})$ then $T^{\alpha}:=T_{1}^{\alpha_{1}}\cdots T_{n}^{\alpha_{n}}$. The coefficient $c_{\alpha}$ is the number of intersection points of $Z$ with a general linear space of the form
$\mathcal{L}_{\alpha}:=\cap_{i=1}^{n}V(\ell_{i,1},\dots,\ell_{i,\alpha_{i}})$
where $\ell_{i,j}$ is a general linear form in the unknowns $\overline{\mathbf{x}}_{i}$.
If $Z$ is reducible, then the multidegree of $Z$ is defined to be the sum of the multidegrees of the irreducible components. If $Z_{1}$, $Z_{2}$ are each equidimensional and intersect transversally, then
$\operatorname{Deg}(Z_{1}\cap Z_{2})=\operatorname{Deg}(Z_{1})\cdot\operatorname{Deg}(Z_{2}).$
Example 1.1.
Consider the bi-homogeneous polynomials
$f_{d}=x_{0}y_{0}^{d}+x_{1}y_{1}^{d}+x_{2}y_{2}^{d}+x_{3}y_{3}^{d},\quad d=0,1,2,3.$
Each polynomial $f_{d}$ defines a hypersurface in $\mathbb{P}^{3}\times\mathbb{P}^{3}$. The multidegree of $V(f_{d})$ is $T_{1}+dT_{2}.$ The multidegree of $V(f_{1},f_{2},f_{3})$ is
$(T_{1}+1T_{2})(T_{1}+2T_{2})(T_{1}+3T_{2})=T_{1}^{3}+6T_{1}^{2}T_{2}+11T_{1}T_{2}^{2}+6T_{2}^{3}$
and the multidegree of $V(f_{1},f_{2},f_{3})\cap V(f_{0})$ is $6T_{1}^{3}T_{2}+11T_{1}^{2}T_{2}^{2}+6T_{1}T_{2}^{3}.$ $\diamond$
1.2 Multiprojective versus affine multiview varieties
In multiview geometry, the natural setting is multiprojective space. However, optimization problems such as Euclidean distance minimization are typically formulated in an affine chart. We now describe both perspectives to get a multiprojective and an affine multiview variety.
An arrangement $\mathbf{C}=(C_{1},\dots,C_{n})$ of full rank $3\times 4$ matrices, called cameras, gives a multiprojective multiview variety of $\mathbb{P}^{3}$ as the Zariski closure of the image of the rational map
$\Phi_{\mathbf{C}}:\mathbb{P}^{3}$ $\dashrightarrow(\mathbb{P}^{2})^{n}$ (1.1)
$z$ $\mapsto(C_{1}z,\dots,C_{n}z).$
As in Section 1.1, we take the coordinates of $(\mathbb{P}^{2})^{n}$ to be $\overline{\mathbf{x}}_{1},\dots,\overline{\mathbf{x}}_{n}$ The multidegree of the image of (1.1) is
$\sum_{\max\{\alpha_{1},\dots,\alpha_{n}\}=2,\,\,\alpha_{1}+\cdots+\alpha_{n}=3}1\cdot{T_{1}^{2-\alpha_{1}}\cdot T_{2}^{2-\alpha_{2}}\cdot\cdots\cdot T_{n}^{2-\alpha_{n}}},\quad n>1.$ (1.2)
For a textbook introduction to multidegrees see [23, Chapter 8]. Ideals of multiview varieties are also well studied in algebraic vision, and for more details we highlight [1, 3, 9].
The affine multiview variety of $\mathbb{P}^{3}$ is obtained by restricting each factor of $(\mathbb{P}^{2})^{n}$ to an affine chart. Our convention is to take the charts given by $x_{i,0}=1$ for $i=1,\dots,n$. Under this choice, the affine multiview variety has the following parameterization:
$z\mapsto\begin{bmatrix}\displaystyle{\frac{C_{1}^{(1)}\cdot z}{C_{1}^{(0)}\cdot z}}&\dots&\displaystyle{\frac{C_{n}^{(1)}\cdot z}{C_{n}^{(0)}\cdot z}}\\ \\ \displaystyle{\frac{C_{1}^{(2)}\cdot z}{C_{1}^{(0)}\cdot z}}&\dots&\displaystyle{\frac{C_{n}^{(2)}\cdot z}{C_{n}^{(0)}\cdot z}}\end{bmatrix}\in\mathbb{K}^{2\times n}.$
In general, if $Z$ is a multiprojective variety then we denote its restriction to the conventional affine chart by $Z_{{\operatorname{Aff}}}\subset(\mathbb{K}^{h})^{n}$.
1.3 Euclidean distance degree
The Euclidean distance degree (ED degree) of an algebraic variety measures the algebraic complexity of the problem of minimizing Euclidean distance to that variety. This invariant plays a central role in applications such as computer vision, statistics, and optimization. We recall its definition in this subsection.
If $X$ is an affine variety in ${{\mathbb{C}}}^{h}$, then we define $X_{\mathbb{R}}:=X\cap\mathbb{R}^{h}$ to be the real points of $X$. We call $u\in\mathbb{R}^{h}$ a data point. The optimization problem
$\min_{x\in X_{\mathbb{R}}}\|x-u\|^{2}$
has first-order optimality conditions. Namely, for a smooth point $x\in X_{\mathbb{R}}$, we require $(x-u)^{T}v=0$ for all $v$ in the tangent space $T_{x}X_{\mathbb{R}}$ of $X_{\mathbb{R}}$ at $x$. This condition gives rise to a polynomial system [6, Equation 2.1] that has $x$ as the unknown and $u$ as the given generic data. The number of complex (real or nonreal) solutions to this polynomial system is called the Euclidean distance degree of $X_{\mathbb{R}}$ [6]. The ED degree is a measure of complexity for minimizing the Euclidean distance from $X_{\mathbb{R}}$ to $u$.
Example 1.2.
Let $X$ denote the variety of the ideal generated by
$g_{1}=x^{2}+y^{2}+z^{2}-1,\quad g_{2}=x^{4}+y^{4}+z^{4}+ax^{2}y^{2}+by^{2}z^{2}+cz^{2}x^{2}-s\left(x^{2}+y^{2}+z^{2}\right)^{2}$
where (a,b,c,s)=(0.8,0.74,0.5,0.32)(a,b,c,s)=(-0.8,-0.74,-0.5,0.32). Then, for a generic data point $u$, the ideal of the critical points is generated by $g_{1},g_{2}$ and the determinant of
$\begin{bmatrix}\nabla d_{u}&\nabla g_{1}&\nabla g_{2}\end{bmatrix}$
where $d_{u}$ is the squared Euclidean distance function from the data $u$ and $\nabla g$ denotes the $3\times 1$ matrix of partial derivatives of $g$. In Figure 2, the silver curve is the restriction of the variety of the determinant to the sphere and the real ED critical points are the intersection points of the silver and red curves.
Note, when $X$ is singular an additional step of saturation needs to be performed to obtain the ideal of critical points. $\diamond$
Figure 2: The red curve is $X_{\mathbb{R}}$ and the real ED critical points for the distance function to the orange points is the set of intersection points with the silver curve.
Beyond its algebraic definition, the Euclidean distance degree also has a topological interpretation. For more details, we refer to the survey article [22].
Theorem 1.3 ([20]).
Suppose $X$ is an irreducible closed subvariety of $\mathbb{C}^{h}$. Then for general $\boldsymbol{\beta}:=(\beta_{0},\dots,\beta_{h})\in\mathbb{C}^{h+1}$ we have
$\text{EDDeg}(X)=(-1)^{\dim X}\chi(\operatorname{Eu}_{X}|_{X\cap U_{\boldsymbol{\beta}}})$
where $\operatorname{Eu}_{X}$ is the local Euler obstruction function on $X$ and
$U_{\boldsymbol{\beta}}:=\mathbb{C}^{h}\setminus\{(z_{1},\dots,z_{h})\in\mathbb{C}^{h}:\sum_{i=1}^{h}(z_{i}-\beta_{i})^{2}+\beta_{0}=0\}.$
For many of our purposes, $X$ is smooth and it suffices to consider the Euler-Poincare characteristic $\chi$.
Theorem 1.4 ([20]).
Suppose $X$ is a smooth closed subvariety of $\mathbb{C}^{h}$. Then for general $\boldsymbol{\beta}:=(\beta_{0},\dots,\beta_{h})\in\mathbb{C}^{h+1}$ we have
$\text{EDDeg}(X)=(-1)^{\dim X}\chi(X\cap U_{\boldsymbol{\beta}})$
where $U_{\boldsymbol{\beta}}:=\mathbb{C}^{h}\setminus\{(z_{1},\dots,z_{h})\in\mathbb{C}^{h}:\sum_{i=1}^{h}(z_{i}-\beta_{i})^{2}+\beta_{0}=0\}$.
In Section 2 we will apply these results to compute the ED degree of one dimensional multiview varieties.
1.4 The projective ED degree
The projective ED degree provides a natural way to study ED degrees of affine cones by avoiding the singularity at the cone point and enabling the use of intersection-theoretic tools. In this subsection, we recall this invariant and prior work for curves.
Let $Z\subset\mathbb{P}^{h}$ be an irreducible closed subvariety of codimension $c$ and let $I_{Z}$ be its defining ideal. If the homogeneous polynomials $g_{1},\dots,g_{s}\in\mathbb{C}[z_{0},z_{1},\dots,z_{h}]$ generate $I_{Z}$, then the projective Euclidean distance degree correspondence is a biprojective variety in $\mathbb{P}^{h}\times\mathbb{P}^{h}$ defined by the saturated ideal
$\biggl(I_{Z}+\biggl\langle\hbox{$(c+2)\times(c+2)$-minors of $\begin{pmatrix}u\\ z\\ J(g)\end{pmatrix}$}\biggr\rangle\biggr):\bigl(I_{Z_{\rm sing}}\cdot\langle z_{0}^{2}+\cdots+z_{h}^{2}\rangle\,\bigr)^{\infty}$ (1.3)
where $J(g)$ is the $s\times(h+1)$ matrix of partial derivatives of $g$ and the singular locus $Z_{\operatorname{sing}}$ of $Z$ is given by the ideal
$I_{Z_{\operatorname{sing}}}\,\,=\,\,I_{Z}+\bigl\langle\hbox{$(c\times c)$-minors of $J(g)$}\bigr\rangle.$
The ED bidegree is defined to be multidegree of the ED degree correspondence. The coefficient of $T_{2}^{h}$ is the (projective) ED degree of $X$, and is denoted by $\operatorname{EDdeg_{proj}}(X)$. The other coefficients of the multidegree are known as the sectional ED degree, and the degree of $X$ is the coefficient of $T_{1}^{c}T_{2}^{h-c}$.
The affine cone over $Z$ is
$\{(z_{0},\dots,z_{h})\in\mathbb{C}^{h+1}:[z_{0}:\dots:z_{h}]\in Z\}.$
The ED degree of this affine cone agrees with $\operatorname{EDdeg_{proj}}(Z)$ [6]. The advantage of working projectively is to avoid the singularity at the cone point. It also allows the use of several topological formulas.
Theorem 1.5 (Theorem 1.3 [21]).
Let $Z$ be any irreducible closed subvariety of $\mathbb{P}^{h}$. Then for a general ${{\boldsymbol{\beta}}}=(\beta_{0},\ldots,\beta_{h})\in{{\mathbb{C}}}^{{h+1}}\setminus\{0\}$ we have
$\operatorname{EDdeg_{proj}}(Z)=(-1)^{\dim X}\chi\big({\rm Eu}_{Z}|_{\mathcal{U}_{{\boldsymbol{\beta}}}}\big),$
where
$\mathcal{U}_{{\boldsymbol{\beta}}}:=\mathbb{P}^{{h}}\setminus(Q\cup H_{{{\boldsymbol{\beta}}}})$
with $Q$ denoting the quadric $\{[z_{0}:\dots:z_{h}]\in\mathbb{P}^{{h}}:\sum_{i=0}^{h}z_{i}^{2}=0\}$ and
$H_{{{\boldsymbol{\beta}}}}:=\{[z_{0}:\dots:z_{h}]\in\mathbb{P}^{{h}}:\beta_{0}x_{0}+\ldots+\beta_{{h}}x_{h}=0\}.$
When $Z\subset\mathbb{P}^{h}$ is smooth, the formula simplifies.
Theorem 1.6.
[2, Theorem 8.1] Let $Z$ be a smooth subvariety of $\mathbb{P}^{h}$, and assume that $Z\nsubseteq Q$, where $Q=\{[z_{0}:\ldots:z_{h}]\in\mathbb{P}^{h}\mid z_{0}^{2}+\ldots+z_{h}^{2}=0\}$ is the quadric in $\mathbb{P}^{h}$. Then
$\operatorname{EDdeg_{proj}}(Z)=(-1)^{\dim Z}\chi\big(Z\setminus(Q\cup H)\big),$
where $H$ is a general hyperplane.
As noted in [2, Section 9], an immediate application of the previous theorems gives a formula for the ED degree of a smooth curve $Z$ in $\mathbb{P}^{h}$ of degree $d$:
$\operatorname{EDdeg_{proj}}(Z)=d+\#(Q\cap Z)-\chi(Z).$ (1.4)
The next examples illustrate some of the discrepancies that occur between an affine variety’s ED degree (which is our setting) and the ED degree of related projective varieties.
Example 1.7 (Standard grading giving discrepancy).
Consider the curve $Z$ in $\mathbb{P}^{2}\times\mathbb{P}^{2}$ parameterized by
$\mathbb{P}^{1}\to Z\subset\mathbb{P}^{2}\times\mathbb{P}^{2}$
[…]
The affine variety $Z_{\operatorname{Aff}}\subset\mathbb{K}^{4}$ has ED degree 10. On the other hand, we can regard the ideal of $Z$ as a homogeneous ideal in the usual $\mathbb{Z}$-grading. This defines a projective surface in $\mathbb{P}^{5}$ that has ED degree 14. $\diamond$
Example 1.8 (Projective closure discrepancy).
The affine $n$-view variety discussed in Section 1.2 is known [20] to have ED degree
$\frac{9}{2}n^{3}-\frac{21}{2}n^{2}+8n-4.$
However, by [12, Theorem 1], its projective closure in $\mathbb{P}^{2n}$ has ED degree
$6n^{3}-15n^{2}+11n-4.\vskip-24.0pt$
$\diamond$
1.5 More multiview varieties and triangulation
In this subsection, we define what are known as anchored point multiview varieties. To stay organized and minimize notation, we focus on the camera arrangement and the image it induces.
Definition 1.9.
Let $\mathbf{C}=(C_{1},\dots,C_{n})$ be a camera arrangement, where each $C_{i}$ is a full-rank $(h+1)\times(N+1)$ matrix. Define the rational map
$\Phi_{\mathbf{C}}:\mathbb{P}^{N}\dashrightarrow(\mathbb{P}^{h})^{n},\quad X\mapsto(C_{1}X,\dots,C_{n}X).$ (1.5)
The (point) multiview variety of $\mathbb{P}^{N}$ with respect to $\mathbf{C}$ is
$\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\mathbb{P}^{N}:=\overline{\operatorname{im}(\Phi_{\mathbf{C}})}\subset(\mathbb{P}^{h})^{n},$
where the overline denotes the Zariski closure. If $Y\subseteq\mathbb{P}^{N}$ is a subvariety, we define
$\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y:=\overline{\Phi_{\mathbf{C}}(Y)}$
and call this the (point) multiview variety anchored at $Y$ with respect to $\mathbf{C}$. We denote its restriction to the standard affine chart by $(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y)_{{\operatorname{Aff}}}$ and refer to it as the affine multiview variety of $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y$.
When $N=3$ and $h=2$, the map $\Phi_{\mathbf{C}}$ models the procedure of taking $n$ pictures with $n$ possibly distinct cameras. A point $P\in\mathbb{P}^{3}$ is mapped to the tuple of images $(C_{1}P,\dots,C_{n}P)$ provided that $P$ is not in the kernel of any $C_{i}$. Points $P\in\mathbb{P}^{N}$ that lie in the kernel of some camera are referred to as being at a camera center, where $\Phi_{\mathbf{C}}$ is undefined. That is, the camera centers constitute the base locus of $\Phi_{\mathbf{C}}$.
Example 1.10 (A one dimensional multiview variety).
If $Y$ is a line in $\mathbb{P}^{3}$, consider the anchored multiview variety $(C_{1},\dots,C_{n})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y$. The affine multiview variety of $(C_{1},\dots,C_{n})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y$ has ED Degree 3n23n-2. This result appears in [24, Theorem 1.7] and serves as a baseline case for understanding the complexity of triangulating points on curves. $\diamond$
In [7], they avoid the subscript $\operatorname{Aff}$ by defining the ED degree of $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\mathbb{P}^{N}$ to be the ED degree of $(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\mathbb{P}^{N})_{\operatorname{Aff}}$. They mention how this choice is inconsistent with [6] in the one view setting. To avoid confusion, we opt to use the following notation. If $X$ is a multiprojective variety, then
$\operatorname{affEDdeg}(X):=\operatorname{EDdeg}(X_{\operatorname{Aff}})$ (1.6)
where $X_{\operatorname{Aff}}$ is the restriction of $X$ to the conventional affine chart described in Section 1.2.
1.6 A Schubert variety in $\operatorname{Gr}(1,\mathbb{P}^{3})$
In this subsection, we define a subvariety of the Grassmannian that will induce the multiprojective varieties appearing in our results in Section 3. Let $\operatorname{Gr}(1,\mathbb{P}^{3})$ denote the Grassmannian of lines in $\mathbb{P}^{3}$. Given three skew lines ${\ell^{(1)}},{\ell^{(2)}},{\ell^{(3)}}$ we define
${\texttt{L}^{\texttt{3}}}({\ell^{(1)}},{\ell^{(2)}},{\ell^{(3)}}):=\{{\ell}\in\operatorname{Gr}(1,\mathbb{P}^{3}):{\ell}\cap{\ell^{(i)}}\neq\emptyset,\,i=1,2,3\}.$ (1.7)
This is an example of a Schubert variety [15]. We often abbreviate ${\texttt{L}^{\texttt{3}}}({\ell^{(1)}},{\ell^{(2)}},{\ell^{(3)}})$ by $\texttt{L}^{\texttt{3}}$}. To understand the geometry of $\texttt{L}^{\texttt{3}}$}, we recall two classical results about lines and quadrics in $\mathbb{P}^{3}.$
Theorem 1.11.
There is a unique smooth quadric $S$ that contains any three given pairwise disjoint lines in $\mathbb{P}^{3}$.
A consequence of Theorem 1.11 is that any line ${\ell}\in\operatorname{Gr}(1,\mathbb{P}^{3})$ satisfying the incidence condition (1.7) must intersect the quadric at three points and is therefore contained in it. Thus, $\texttt{L}^{\texttt{3}}$} lies within the family of lines contained in the quadric. The next theorem characterizes the lines contained in a quadric.
Theorem 1.12.
A smooth quadric $S$ in $\mathbb{P}^{3}$ contains two rulings. Moreover, any two lines from different rulings intersect in exactly one point, and any two lines from the same ruling are disjoint.
Theorem 1.12 implies that the three skew lines in our incidence condition are in the same ruling. All together, these theorems imply that $\texttt{L}^{\texttt{3}}$} is the other ruling.
The Plücker embedding
$\iota:\operatorname{Gr}(1,\mathbb{P}^{3})\to\mathbb{P}^{5},$
provides us a way to embed ${\texttt{L}^{\texttt{3}}}\subset\operatorname{Gr}(1,\mathbb{P}^{3})$ into $\mathbb{P}^{5}$. It turns out $\iota({\texttt{L}^{\texttt{3}}})$ is a conic curve in $\mathbb{P}^{5}$, and the next example illustrates how to parameterize it.
Example 1.13 (Parameterizing $\texttt{L}^{\texttt{3}}$}).
In this example, we describe the parametrization of $\iota({\texttt{L}^{\texttt{3}}})$ where $\ell^{(i)}$} is the line
$\mathbb{P}^{1}\to\mathbb{P}^{3},\,\quad[s:t]\mapsto[su_{i}:sv_{i}:tu_{i}:tv_{i}]\quad\text{ with fixed }\quad[u_{i}:v_{i}]\in\mathbb{P}^{1}.$ (1.8)
The unique smooth quadric $S$ containing these lines is the image of the Segre embedding
$\sigma:\mathbb{P}^{1}\times\mathbb{P}^{1}$ $\to\mathbb{P}^{3}$
$([s:t],[u:v])$ $\mapsto[su:sv:tu:tv]$
so that $S$ is defined by the equation $x_{0}x_{3}-x_{1}x_{2}=0$. The line $\ell^{(i)}$} is spanned by the points $u_{i}:v_{i}:0:0]$ and $0:0:u_{i}:v_{i}]$, and we get a ruling in $S$ containing the three skew lines by varying the point $u:v]\in\mathbb{P}^{1}$. We get the second ruling, $\texttt{L}^{\texttt{3}}$}, by varying $s:t]\in\mathbb{P}^{1}$ to get lines spanned by the points $[s:0:t:0]$ and $[0:s:0:t]$.
In terms of Plücker coordinates on $\mathbb{P}^{5}$, lines in $\iota({\texttt{L}^{\texttt{3}}})$ are given by
$\begin{bmatrix}p_{12}\\ p_{13}\\ p_{23}\\ p_{14}\\ p_{24}\\ p_{34}\end{bmatrix}=\begin{bmatrix}s^{2}\\ 0\\ -st\\ st\\ 0\\ t^{2}\end{bmatrix}$ (1.9)
To parameterize the lines in $\iota({\texttt{L}^{\texttt{3}}})$ by $\mathbb{P}^{1}$, we use a composition of the Veronese embedding
$\nu([s:t])=\begin{bmatrix}s^{2}\\ st\\ t^{2}\end{bmatrix}$
with a linear map $B_{\texttt{L}^{\texttt{3}}}$ represented by the matrix
$B_{{\texttt{L}^{\texttt{3}}}}=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&-1&0\\ 0&1&0\\ 0&0&0\\ 0&0&1\end{bmatrix}$
so that the parameterization map $f:\mathbb{P}^{1}\to\iota({\texttt{L}^{\texttt{3}}})$ takes $[s:t]$ to $B_{{\texttt{L}^{\texttt{3}}}}\cdot\nu([s:t])$.
Note, our choice for the three skew lines ${\ell^{(1)}},{\ell^{(2)}},{\ell^{(3)}}$ in (1.8) was made without loss of generality as the unique quadric containing any other choice of three skew lines is projectively equivalent to $V(x_{0}x_{3}-x_{1}x_{2})$. $\diamond$
Remark 1.14.
If $\mathbf{D}$ is a camera arrangement of size $(H+1)\times 6$, then $\mathbf{D}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\iota({\texttt{L}^{\texttt{3}}})\subset(\mathbb{P}^{H})^{n}$ is a point multiview variety anchored at $\iota({\texttt{L}^{\texttt{3}}})$. This variety was previously studied in [7] when $\mathbf{D}$ is an arrangement of so-called wedge cameras (Section 3.2). We revisit this construction and show how it relates to line multiview varieties in Section 3 for $H=\binom{2+1}{2}-1$ and $H=\binom{3+1}{2}-1$.
2 ED degree of multiview varieties of a rational curve
In this section, we let $Y$ be a degree $e$ rational curve in $\mathbb{P}^{N}$ with $N\geq 3$ parametrized by
$f:\mathbb{P}^{1}$ $\to\mathbb{P}^{N}$ (2.1)
$[s:t]$ $\mapsto f(s,t)=[f_{0}(s,t):f_{1}(s,t):\dots:f_{N}(s,t)].$
We say $Y$ is the world variety. We always assume the $f_{i}$ are homogeneous degree $e$ polynomials such that $f$ has no base locus and is generically one-to-one. Unless otherwise stated, we also assume $Y=f(\mathbb{P}^{1})$ is smooth or only has nodes as singularities.
As a special case of Section 1.5, for $Y$ as in (2.1) we denote its multiview variety with respect to the camera arrangement $\mathbf{C}$ as $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})$. If $Y$ does not contain any points in a camera center of the arrangement, then the image $\Phi_{\mathbf{C}}(f(\mathbb{P}^{1}))$ is closed.
Lemma 2.1.
Assume we are given a generic camera arrangement $\mathbf{C}=(C_{1},\dots,C_{n})$ of full rank $(h+1)\times(N+1)$ matrices $C_{i}$ with $N\geq 3$ and $h\geq 2$. Let $f:\mathbb{P}^{1}\to\mathbb{P}^{N}$ be as in (2.1) so that $Y=f(\mathbb{P}^{1})$ is degree $e$ and $Y$ is either smooth or has only nodal singularities.
Then $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})$ is either a smooth curve or only has nodes as singularities.
Proof.
For $n=1$, this is a standard algebraic geometry fact. For instance, see [8, Exercise 3.34]. The case when $n>1$ immediately follows. ∎
Example 2.2 (Single view varieties of cubic curves).
Let $Y=f(\mathbb{P}^{1})$ be the rational cubic curve in $\mathbb{P}^{3}$ of degree $3$ given by
$s:t]\overset{f}{\mapsto}[t^{3}:st^{2}:s^{2}t:s^{3}].$
Let $C$ be a general $(h+1)\times{4}$ camera matrix and take the camera configuration to be this single camera ($n=1$).
There are three different cases:
  1. 1.
    If $h=1$, then $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})$ is $\mathbb{P}^{1}$. The ED degree of $(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1}))_{\operatorname{Aff}}$ is one.
  2. 2.
    If $h=2$, then $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})$ is a curve in $\mathbb{P}^{2}$ with a nodal singularity.
  3. 3.
    If $h\geq 3$, then $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})$ is a smooth curve in $\mathbb{P}^{h}$.
If $h\geq 2$ then the ED degree of $(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1}))_{\operatorname{Aff}}$ is seven. In Figure 1 (left) we plot $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})$ in orange for a camera of size $3\times 4$ and the curve $Y$ in yellow.
$\diamond$
Theorem 2.3.
Assume we are given a generic camera arrangement $\mathbf{C}=(C_{1},\dots,C_{n})$ of full rank $(h+1)\times(N+1)$ matrices $C_{i}$ with $N\geq 3$ and $h\geq 2$. Let $f:\mathbb{P}^{1}\to\mathbb{P}^{N}$ be as in (2.1) so that $Y=f(\mathbb{P}^{1})$ is degree $e$ and $Y$ is smooth or has only nodal singularities. Then the ED degree of an affine patch $(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1}))_{\operatorname{Aff}}$ of $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})$ is
$\operatorname{affEDdeg}(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1}))=3en-2.$ (2.2)
Proof.
Recall from Section 1.5 that a camera $C$ is a full rank $(h+1)\times(N+1)$ matrix. We write $C^{(i)}$ for the $i$th row of $C$. The subvariety $X:=\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})$ has the parameterization
$\Phi_{\mathbf{C}}\circ f([s:t])=\begin{bmatrix}\begin{pmatrix}C_{1}^{(0)}\cdot f\\ C_{1}^{(1)}\cdot f\\ \vdots\\ C_{1}^{(h)}\cdot f\end{pmatrix}:\dots:\begin{pmatrix}C_{n}^{(0)}\cdot f\\ C_{n}^{(1)}\cdot f\\ \vdots\\ C_{n}^{(h)}\cdot f\end{pmatrix}\end{bmatrix}\in(\mathbb{P}^{h})^{n}.$ (2.3)
Let $X_{\operatorname{Aff}}:=(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y)_{\operatorname{Aff}}$ denote the restriction of $\operatorname{im}\Phi_{\mathbf{C}}$ to an affine patch with coordinates
$\begin{pmatrix}\begin{pmatrix}\frac{C_{1}^{(1)}\cdot f}{C_{1}^{(0)}\cdot f}\\ \vdots\\ \frac{C_{1}^{(h)}\cdot f}{C_{1}^{(0)}\cdot f}\end{pmatrix},\dots,\begin{pmatrix}\frac{C_{n}^{(1)}\cdot f}{C_{n}^{(0)}\cdot f}\\ \vdots\\ \frac{C_{n}^{(h)}\cdot f}{C_{n}^{(0)}\cdot f}\end{pmatrix}\end{pmatrix}\in(\mathbb{K}^{h})^{n}.$ (2.4)
In $(\mathbb{P}^{h})^{n}$, each factor of $\mathbb{P}^{h}$ gives a hyperplane at infinity. We denote the union of these hyperplanes as $H_{\infty}$. If the $i$th factor in $\mathbb{P}^{h}$ has coordinates $\overline{\mathbf{x}}_{i}=(x_{i,0},x_{i,1},\dots,x_{i,h})$ then
$H_{\infty}=\cup_{i=1}^{n}V(x_{i,0}).$
Using the parameterization of $X$ in (2.3), we have
$X\cap H_{\infty}=\Phi_{\mathbf{C}}\circ\phi(S_{\infty})\text{\quad where\quad}S_{\infty}:=\bigcup_{i=1}^{n}V(C_{i}^{(0)}\cdot f)\subset\mathbb{P}^{1}.$
If each $C_{i}^{(0)}\cdot f$ has distinct roots and
$V(C_{i}^{(0)}\cdot f,C_{j}^{(0)}\cdot f)=\emptyset$ (2.5)
for all $i,j\in[n],\ i\neq j$, then $S_{\infty}$ has $ne$ points. This holds by genericity of the camera arrangement: the degree $e$ polynomial $C_{i}^{(0)}\cdot f$ has a multiple root if and only if its $X$-discriminant vanishes, and $C_{i}^{(0)}\cdot f,C_{j}^{(0)}\cdot f$ share a root if and only if their resultant is nonvanishing [11, Section 1.1]. Therefore, if $f$ is a generically one-to-one map then $H_{\infty}\cap X$ has precisely $ne$ points.
For $\beta=(\beta_{0},\beta_{1,1},\dots,\beta_{n,h})\in\mathbb{R}^{nh+1}$, define the hypersurface
$Q_{\beta}:=\left\{(z_{1,1},\dots,z_{n,h})\in\mathbb{C}^{nh}:\sum_{i=1}^{n}\sum_{j=1}^{h}(z_{i,j}-\beta_{i,j})^{2}+\beta_{0}=0\right\}$ (2.6)
and the Zariski open set
$U_{\beta}:=\mathbb{C}^{hn}\setminus Q_{\beta}.$ (2.7)
Let $S_{Q}$ be the set of points in $\mathbb{P}^{1}\setminus S_{\infty}$ that are the zeros of
$d_{\beta}=\sum_{i=1}^{n}\sum_{j=1}^{h}\left(\frac{C_{i}^{(j)}\cdot f}{C_{i}^{(0)}\cdot f}-\beta_{i,j}\right)^{2}+\beta_{0}.$
By construction, we have $(\Phi_{\mathbf{C}}\circ\phi(S_{Q}))_{\operatorname{Aff}}=X_{\operatorname{Aff}}\cap Q_{\beta}$. We may write $d_{\beta}$ as a quotient of polynomials
$d_{\beta}=\frac{g_{\beta}}{D}\quad\text{ with }\quad D:={\prod_{i=1}^{n}\left(C_{i}^{(0)}\cdot f\right)^{2}}$
where
$g_{\beta}=\beta_{0}D+\sum_{i=1}^{n}\frac{D}{\left(C_{i}^{(0)}\cdot f\right)^{2}}\sum_{j=1}^{h}\left({(C_{i}^{(1)}\cdot f)}-\beta_{i,j}{(C_{i}^{(0)}\cdot f})\right)^{2}$
is a homogeneous degree $2ne$ polynomial in $s,t$. Because $\beta_{0}$ is generic, if $g_{\beta}$ has no common roots with $D$ then $g_{\beta}$ has $2ne$ distinct roots. On the other hand, if $g_{\beta}$ had a common root with $D$, then there would exist a unique $i=a$ such that $\left(C_{i}^{(0)}\cdot f\right)$ vanishes on that root. This implies the polynomial
$\sum_{j=1}^{h}{\left(C_{a}^{(j)}\cdot f\right)}^{2}$
vanishes at the root. By the genericity of $C_{a}^{(j)}\cdot f$ this cannot happen, thus $g_{\beta}$ has $2en$ roots.
The degree of $g_{\beta}$ equals $\chi(X\cap Q_{\beta}).$ It can be obtained by counting the degree of, for instance, the first term as we clear denominators. Multiplying through by $C_{1}^{(0)}\cdot f^{2}$ gives a leading monomial of degree $2en$, to which multiplication with subsequent $C_{i}^{(0)}\cdot f^{2}$ adds $2en$ for each pair of terms from the same camera. Thus, in an $n$-camera configuration, $\chi(X_{\operatorname{Aff}}\cap Q_{\beta})=2en$.
If $X_{\operatorname{Aff}}$ is smooth, an application of [20, Theorem 3.8] implies that
$\operatorname{affEDdeg}(X)=-(\chi(X)-\chi(X\cap H_{\infty})-\chi(X\cap Q_{\beta})).$ (2.8)
Hence $\operatorname{affEDdeg}(X)=-(\chi(\Phi_{\mathbf{C}}\circ\phi(\mathbb{P}^{1}))-\chi(\Phi_{\mathbf{C}}\circ\phi(S_{\infty}))-\chi(\Phi_{\mathbf{C}}\circ\phi(S_{Q})))$, and therefore
$\operatorname{affEDdeg}(X)=(-1)(2-en-2en)=3en-2.$ (2.9)
If $X_{\operatorname{Aff}}$ is singular then we use Theorem 1.3, which implies for general $\boldsymbol{\beta}:=(\beta_{0},\beta_{1},\dots,\beta_{hn})\in\mathbb{C}^{hn+1}$ that $\operatorname{affEDdeg}(X)$ equals
$\left(\chi(X_{{\operatorname{reg}}})-\chi(X_{{\operatorname{reg}}}\cap H_{\infty})-\chi((X_{\operatorname{reg}})_{\operatorname{Aff}}\cap Q_{\beta})+\sum_{x\in\text{Sing}(X_{\operatorname{Aff}})\setminus Q_{\beta}}m_{x}\right)$
where $X_{\operatorname{reg}}$ denotes the smooth points of $X$ and $m_{x}$ denotes the multiplicity of $x$. Similar arguments as in the smooth case show $\chi(X_{\operatorname{reg}}\cap H_{\infty})=en$ and
$\chi((X_{\operatorname{reg}})_{\operatorname{Aff}}\cap Q_{\beta})=2en$
as a consequence of the genericity of $\mathbf{C}$. Also by the genericity of $\mathbf{C}$ we have
$\sum_{x\in\text{Sing}(X_{\operatorname{Aff}})\setminus Q_{\beta}}m_{x}=\sum_{x\in\text{Sing}(X)}m_{x}.$
Substituting in the known quantities, we find
$\operatorname{affEDdeg}(X)=(-1)\left(\chi(X_{{\operatorname{reg}}})-en-2en+\sum_{x\in\text{Sing}(X)}m_{x}\right).$ (2.10)
Since each point in the singular locus of $X$ is a node and $X=\Phi_{\mathbf{C}}\circ\phi(\mathbb{P}^{1})$, we have
$\chi(X_{\operatorname{reg}})=\chi(\mathbb{P}^{1})-\sum_{x\in\text{Sing}(X)}m_{x}.$
Plugging into (2.10), we get the desired formula.
By Section 2, the result follows. ∎
Corollary 2.4.
Suppose $f$ is as in (2.1) with $N\geq 3$ and $h\geq 2$. Let $\mathcal{C}\subset\mathbb{P}^{(h+1)\times(N+1)-1}$ be an irreducible projective variety. If $C_{1},\dots,C_{n}$ are generic points in $\mathcal{C}$ and
$\operatorname{affEDdeg}((C_{1},\dots,C_{n})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1}))=3en-2$ (2.11)
holds for $n=1,2$, then (2.11) holds for all $n\geq 1$.
Proof.
When $(C_{1},\dots,C_{n})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})$ is smooth this follows from the proofs of Section 2 and Theorem 2.3, as each genericity condition only involves a single camera or a pair of cameras. ∎
Example 2.5 (A non-generic family).
Let $Y$ be the rational curve in $\mathbb{P}^{5}$ of degree $5$ given by
$s:t]\overset{f}{\mapsto}[t^{5}:st^{5-1}:\dots:s^{5}].$
Define the following irreducible variety of cameras:
$\mathcal{C}:=\left\{\begin{bmatrix}c_{1}&c_{2}&0&0&0&0\\ 0&0&c_{3}&c_{4}&0&0\\ 0&0&0&0&c_{5}&c_{6}\end{bmatrix}\in\mathbb{K}^{3\times 6}:c_{i}\in\mathbb{K}\right\}.$
If $C$ is a generic element in $\mathcal{C}$, then $\operatorname{affEDdeg}(C_{1}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y)=9$. We see that this disagrees with the numbers in (2.11).
$\diamond$
Example 2.6 (Non-generic two views).
Consider the following family of $3\times 4$ camera matrices:
$\mathcal{C}=\left\{[c_{i,j}]\in\mathbb{K}^{3\times 4}:c_{1,1}=0\right\}.$
Let $C_{1}$ and $C_{2}$ be generic points of $\mathcal{C}$. If $Y$ is the rational normal curve in $\mathbb{P}^{3}$ parameterized by $t^{3}:st^{2}:s^{2}t:s^{3}]$, then
$\operatorname{affEDdeg}(C_{1}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y)=7\text{ and }\operatorname{affEDdeg}((C_{1},C_{2})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y)=13.$
The single view case agrees with the ED degrees given by the formula in Theorem 2.3, but the 2-view case does not. $\diamond$
Example 2.7 (Smooth to singular).
Let $Y$ be the rational curve in $\mathbb{P}^{4}$ of degree $4$ given by
$s:t]\overset{f}{\mapsto}[t^{4}:st^{4-1}:\dots:s^{4}].$
Let $C$ be a general $3\times{(4+1)}$ camera matrix. Then $C\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}X$ is a degree $4$ curve in $\mathbb{P}^{2}$ with $\binom{4-1}{2}$ nodal singularities. The set of smooth points of $C\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}X\subset\mathbb{P}^{2}$ has Euler characteristic $\chi(\mathbb{P}^{1}\setminus\{2\cdot\binom{4-1}{2}\text{ points}\})=2-2\cdot\binom{4-1}{2}$. The ED degree of $(C\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}X)_{\operatorname{Aff}}$ is $3\cdot 4-2$. The projective ED degrees of $X$ and $C\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}X$ both equal $3\cdot 4-2$. As affine cones in ${{\mathbb{C}}}^{4+1}$ and ${{\mathbb{C}}}^{3}$, these varieties are parameterized by degree $4$ polynomials and one might predict from [6, Section 2] that the ED degree is $(2\cdot 4-1)^{2}$. The reason we do not get this bound is because these are general homogeneous polynomials of degree $4$ and not general degree $4$ polynomials. $\diamond$
Example 2.8 (Cuspidal cubic).
If $f(s,t)=[s^{3}:st^{2}:t^{3}]$ and the camera configuration $\mathbf{C}$ consists of a single generic camera, then the ED degree of $(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1}))_{\operatorname{Aff}}$ is six. The cubic curve $f(\mathbb{P}^{1})$ has a cusp as a singularity. Thus, $\chi(X_{\operatorname{reg}})=\chi(\mathbb{P}^{1})-1$ and this is the reason for the drop in the ED degree from the ED degree of a cubic curve in $\mathbb{P}^{2}$ with a node. $\diamond$
Remark 2.9 (Intersection theory and multidegrees).
The results in this section have an intersection theoretic [10] interpretation through the multigrees presented in Section 1.2. Namely, the proof in Theorem 2.3 is showing that the curve $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})\subset(\mathbb{P}^{h})^{n}$ has multidegree $\sum_{i=1}^{n}e\cdot\frac{T_{1}^{h}T_{2}^{h}\cdots T_{n}^{h}}{T_{i}}$ and intersects transversally each of the hyperplanes $V(x_{i,0})$ for $i=1,\dots,n$. This implies $\chi(X\cap H_{\infty})=en$.
In addition, the proof shows that the hypersurface defined by
$q_{\beta}:=\sum_{i=1}^{n}\sum_{j=1}^{h}(x_{i,j}-\beta_{i,j}x_{i,0})^{2}+\beta_{0}\prod_{i=1}^{n}x_{i,0}^{2}$
(obtained by homogenizing the polynomial in (2.6)) also transversally intersects the curve $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})$ with no solutions at infinity. Since the multidegree of $V(q_{\beta})$ is $\sum_{i=1}^{n}2T_{i}^{2}$, we have $\chi(X\cap V(q_{\beta}))=2en$ because
$\left(\sum_{i=1}^{n}2T_{i}^{2}\right)\cdot\left(\sum_{i=1}^{n}e\cdot\frac{T_{1}^{h}\cdots T_{n}^{h}}{T_{i}}\right)=2en\cdot T_{1}^{h}\cdots T_{n}^{h}\in\mathbb{Z}[T_{1},\dots,T_{n}]/\langle T_{1}^{h+1},\dots,T_{n}^{h+1}\rangle.$
3 Strengthening the ED-equivalence catalogue
In this section, we present our main application of our results. Section 3.1 defines our main varieties of interest: a set of multiview varieties anchored at a subvariety of the Grassmannian. Section 3.2 gives a complementary approach for obtaining these multiview varieties using wedge cameras. It also provides the necessary exterior algebra background. Using our results in Section 2, we prove a pair of conjectures in Section 3.3.
3.1 Anchoring the $n$-view variety
Section 1.5 considers the point multiview variety $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y$ anchored at $Y$ as the Zariski closure of the image of the map (1.5) for $Y\subset\mathbb{P}^{N}$. In this section, we carry this idea over to a linear map between Grassmannians. The associated multiview variety (called a generalized multiview variety in [7]) is obtained as the Zariski closure of a rational map
$\Phi_{\mathbf{C},k}:\operatorname{Gr}(k,\mathbb{P}^{N})$ $\dashrightarrow\operatorname{Gr}(k,\mathbb{P}^{h})^{n}$ (3.1)
$P$ $\mapsto(C_{1}\cdot P,\dots,C_{n}\cdot P)$
for a camera arrangement $\mathbf{C}=(C_{1},\dots,C_{n})$ of full-rank $(h+1)\times(N+1)$ matrices. Here, the multiplication $C_{i}\cdot P$ denotes the $k$-plane ($k$-dimensional linear subspace of $\mathbb{P}^{N}$) defined as the span of $C_{i}X_{0},\dots,C_{i}X_{k}$ for $P$ spanned by $X_{0},\dots,X_{k}$. If $k=0$, then $\Phi_{\mathbf{C},k}$ agrees with (1.5).
When the domain of $\Phi_{\mathbf{C},k}$ is restricted to a subvariety, the Zariski closure of the image is still referred to as an anchored multiview variety. Given a subvariety $\Lambda\subseteq\operatorname{Gr}(k,\mathbb{P}^{N})$ of the Grassmannian, one obtains a multiview variety as the Zariski closure of the image of
$\Lambda$ $\dashrightarrow\operatorname{Gr}(k,\mathbb{P}^{h})^{n}$ (3.2)
$P$ $\mapsto(C_{1}\cdot P,\dots,C_{n}\cdot P).$
We summarize this discussion with the following definition.
Definition 3.1.
Let $\Lambda$ be a subvariety of $\operatorname{Gr}(k,\mathbb{P}^{N})$. For a camera arrangement $\mathbf{C}=(C_{1},\dots,C_{n})$ of full-rank $(h+1)\times(N+1)$ matrices, we denote the Zariski closure of the image of (3.2) by $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\Lambda$ and call this the $n$-view variety anchored at $\Lambda$ with respect to $\mathbf{C}$. When $k=1$, we say it is a line $n$-view variety.
Remark 3.2.
In [7], the full data of a multiview variety is recorded in the notation $\mathcal{M}_{\mathbf{C},k}^{\Lambda,N,h}$. We opt for the more parsimonious notation and refer to the catalogue in [7, Theorem 3.3] for a complete description of multiview varieties arising from projections of points and lines in 1, 2, and 3-dimensional projective space.
As defined, $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\Lambda$ is a subvariety of a product of Grassmannians. Next, we embed $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\Lambda$ into a product of projective spaces. This is done using the Plücker embedding on a product of Grassmannians,
$\emph{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}{$\iota_{k,h,n}$}}}:(\operatorname{Gr}(k,\mathbb{P}^{h}))^{n}\to\left(\mathbb{P}^{\binom{h+1}{k+1}-1}\right)^{n}.$ (3.3)
When the context is clear, we suppress the dependence of the map $\iota$ on $k$, $h$, and $n$. Thus $\iota(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\Lambda)_{\operatorname{Aff}}$ is an affine variety contained in $\left({{\mathbb{C}}}^{\binom{h+1}{k+1}-1}\right)^{n}$.
Example 3.3 (Defining $X_{h,n}$).
Recall ${\texttt{L}^{\texttt{3}}}\subset\operatorname{Gr}(1,\mathbb{P}^{3})$ from Section 1.6. Suppose $h=2$ or $h=3$. If $\mathbf{C}^{h}$ is an arrangement of $n$-cameras of size $(h+1)\times 4$, then $\mathbf{C}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}{\texttt{L}^{\texttt{3}}}$ is a subvariety of $\operatorname{Gr}(1,\mathbb{P}^{h})^{n}$. Taking the image of the Plücker embedding
$\iota_{1,h,n}:\mathbf{C}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}{\texttt{L}^{\texttt{3}}}\to(\mathbb{P}^{\binom{h+1}{2}-1})^{n}$
we obtain a multiprojective variety, which we denote by $X_{h,n}$. Table 1 highlights properties of $X_{h,n}$ for small $n$. $\diamond$
Ambient space Multidegree $\operatorname{affEDdeg}(X_{h,n})$
h=2,n=1h=2,\,n=1 $\mathbb{P}^{2}$ $2T_{1}$ 4
h=2,n=2h=2,\,n=2 $\mathbb{P}^{2}\times\mathbb{P}^{2}$ $2T_{1}+2T_{2}$ 10
h=3,n=1h=3,\,n=1 $\mathbb{P}^{5}$ $2T_{1}$ 4
h=3,n=2h=3,\,n=2 $\mathbb{P}^{5}\times\mathbb{P}^{5}$ $2T_{1}+2T_{2}$ 10
Table 1: Summary of $\mathbf{C}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}{\texttt{L}^{\texttt{3}}}$ with $\mathbf{C}^{h}$ having $n=1,2$ cameras in its arrangement.
3.2 Exterior algebra review and wedge cameras
Next, we show how to realize $\mathbf{C}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}{\texttt{L}^{\texttt{3}}}$ as a point multiview variety anchored at $\iota({\texttt{L}^{\texttt{3}}})\subset\mathbb{P}^{5}$. To that end, we review the exterior algebra (see [18, Chapter XVI] for a textbook reference) to state Section 3.2.
Recall the exterior algebra of a vector space $V$ over a field $\mathbb{K}$ is the graded algebra
$\bigwedge V:=\bigoplus_{k=0}^{\dim V}\bigwedge^{k}V,$
where $\bigwedge^{k}V$ is the space of alternating $k$-linear forms on $V$. Elements of $\bigwedge^{k}V$ are called $k$-vectors, and the wedge product satisfies $u\wedge v=-v\wedge u$ and $v\wedge v=0$. A simple wedge is one of the form $v_{1}\wedge\cdots\wedge v_{k}$, $v_{i}\in V.$ Every element of $\bigwedge^{k}V$ is a linear combination of simple wedges. The dimension of $\bigwedge^{k}V$ is $\binom{\dim V}{k}$, and a basis is given by
$e_{i_{1}}\wedge e_{i_{2}}\wedge\cdots\wedge e_{i_{k}},\quad i_{1}<i_{2}<\cdots<i_{k}$
where $\{e_{i}\}$ is a basis for $V$. Any linear map $f:V\to W$ induces the map
$\bigwedge^{k}f:\,\bigwedge^{k}V\longrightarrow\bigwedge^{k}W,\qquad\bigwedge^{k}f(v_{1}\wedge\cdots\wedge v_{k})=f(v_{1})\wedge\cdots\wedge f(v_{k}).$ (3.4)
In our vision setting, each camera matrix $C$ is a linear map $f_{C}:\mathbb{K}^{N+1}\to\mathbb{K}^{h+1}.$ After fixing ordered bases of $\mathbb{K}^{N+1}$ and $\mathbb{K}^{h+1}$, we obtain induced wedge bases on $\bigwedge^{k}\mathbb{K}^{N+1}$ and $\bigwedge^{k}\mathbb{K}^{h+1}$. The wedge camera matrix $\wedge^{k}C$} represents the linear map
$\bigwedge^{k}f_{C}:\;\bigwedge^{k}\mathbb{K}^{N+1}\to\bigwedge^{k}\mathbb{K}^{h+1},\text{ where }v_{1}\wedge\cdots\wedge v_{k}\mapsto f_{C}(v_{1})\wedge\cdots\wedge f_{C}(v_{k}).$
The entries of $\wedge^{k}C$} consist of all $k\times k$ minors of $C$, arranged according to the chosen multi-index ordering.
Definition 3.4.
If $\mathbf{C}=(C_{1},\dots,C_{n})$ is an arrangement of $n$ cameras, then we call $\left({\wedge^{k+1}C_{1}},\dots,{\wedge^{k+1}C_{n}}\right)$ the $(k+1)$-wedge camera arrangement of $\mathbf{C}$.
Example 3.5.
Suppose $C$ is the $3\times 4$ matrix
$C=\begin{bmatrix}1&2&3&4\\ 0&-1&0&0\\ 0&0&5&0\end{bmatrix}$.
Then $f_{\wedge^{2}C}:\;\bigwedge^{2}\mathbb{K}^{4}\to\bigwedge^{2}\mathbb{K}^{3}$ has matrix ${\wedge^{2}C}\in\mathbb{K}^{3\times 6}$ with entries being the $2\times 2$ minors of $C$.
We use the reverse lexicographic ordering $((1,2),(1,3),(2,3))$ on rows and
$((1,2),(1,3),(2,3),(1,4),(2,4),(3,4))$ for columns. Then,
${\wedge^{2}C}=\begin{bmatrix}-1&0&0&3&4&0\\ 0&5&0&10&0&-20\\ 0&0&0&-5&0&0\end{bmatrix}$.
Each entry is $\det C_{I,J}$ with $I$ a row pair and $J$ a column pair in the stated order. $\diamond$
The proof of the following proposition is sketched in [7]. It is a consequence of standard results from multilinear algebra.
Proposition 3.6.
Let $h\geq k$. Fix a camera arrangement $\mathbf{C}=(C_{1},\dots,C_{n})$ of full-rank $(h+1)\times(N+1)$ matrices and a subvariety $\Lambda$ of $\operatorname{Gr}(k,\mathbb{P}^{N})$. Then
$\mathbf{D}=(\wedge^{k+1}C_{1},\dots,\wedge^{k+1}C_{n})$
is a wedge camera arrangement of $\binom{h+1}{k+1}\times\binom{N+1}{k+1}$ full rank matrices. Moreover, $\iota_{k,h,n}(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\Lambda)$ is a point multiview variety anchored at $\iota_{k,N,1}(\Lambda)$:
$\iota_{k,h,n}(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\Lambda)=\mathbf{D}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\iota_{k,N,1}(\Lambda).$
As a consequence of Section 3.2 we get the commutative diagram in Figure 3.
$\Lambda\subset\operatorname{Gr}(k,\mathbb{P}^{N})$$\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\Lambda$$\operatorname{Gr}(k,\mathbb{P}^{h})^{n}$$\iota_{k,N,1}(\Lambda)\subset\mathbb{P}^{\binom{N}{k}}$$\mathbf{D}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\iota_{k,h,n}(\Lambda)$$(\mathbb{P}^{\binom{h+1}{k}-1})^{n}$$\Phi_{\mathbf{C}}$$\iota_{k,N,1}$$\iota_{k,h,n}$$\iota_{k,h,n}$$\Phi_{\mathbf{D}}$}
Figure 3: This commutative diagram shows two approaches for obtaining the $k$-plane multiview variety anchored at $\Lambda$.
3.3 Resolving a pair of ED degree conjectures
Recall the subvariety $\texttt{L}^{\texttt{3}}$} of $\operatorname{Gr}(1,\mathbb{P}^{3})$ from Section 1.6 and Section 3.2. We focus on $\mathbf{C}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}{\texttt{L}^{\texttt{3}}}$, the line view variety anchored at $\texttt{L}^{\texttt{3}}$} with respect to a generic arrangement $\mathbf{C}^{h}$ of $n$ cameras of size $(h+1)\times 4$. Throughout this subsection, we denote the Plücker embedding of $\mathbf{C}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}{\texttt{L}^{\texttt{3}}}$ by
$X_{h,n}:=\iota_{1,h,n}(\mathbf{C}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}{\texttt{L}^{\texttt{3}}}).$
Example 3.7.
Continuing with Section 3.1, suppose $\mathbf{D}^{h}$ is the $2$-wedge camera arrangement of $\mathbf{C}^{h}$. Then $\mathbf{D}^{h}$ is an arrangement of cameras of size $\binom{h+1}{2}\times 6$. By Section 3.2, we have
$X_{h,n}=\mathbf{D}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\iota_{1,3,1}({\texttt{L}^{\texttt{3}}}).$
Furthermore, combining the parameterization in Section 1.6 with the results of Section 3.2 yields the commutative diagram in Figure 4. The commutative diagram shows $\Phi_{\mathbf{D}^{h}}\circ f$ is a parameterization of the line $n$-view variety anchored at $\texttt{L}^{\texttt{3}}$}. $\diamond$
[…]
Figure 4: For $h=2,3$, this shows the equality $\iota_{1,h,n}(\mathbf{C}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}{\texttt{L}^{\texttt{3}}})=\mathbf{D}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\iota_{1,3,1}({\texttt{L}^{\texttt{3}}})$. The unlabeled horizontal arrows are inclusions and $\iota_{k,h,n}$ is as defined in (3.3).
Theorem 3.8 ([7] Conjecture 7.4.5 and Conjecture 7.4.6).
For $h=2,3$, let $\mathbf{C}^{h}$ be a generic configuration of $n$ cameras of size $(h+1)\times 4$. If $X_{n,h}$ denotes $\iota(\mathbf{C}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}{\texttt{L}^{\texttt{3}}})$, then
$\operatorname{affEDdeg}(X_{h,n})$ $=6n-2.$ (3.5)
Proof.
Let $\mathbf{D}^{h}=(D_{1},\dots,D_{n})$ denote the $2$-wedge camera configuration of $\mathbf{C}^{h}$.
By Section 3.2 and Section 3.3, we have
$\iota_{1,h,n}(\mathbf{C}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}{\texttt{L}^{\texttt{3}}})\;$ $=\;\mathbf{D}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\iota_{1,3,1}({\texttt{L}^{\texttt{3}}})$
$X_{h,n}\;$ $=\;\mathbf{D}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\iota_{1,3,1}({\texttt{L}^{\texttt{3}}}).$
Since $\mathbf{C}^{h}$ is generic, each $D_{i}$ is a generic point in the irreducible variety of wedge cameras of size $\binom{h+1}{2}\times 6$. It follows from Figure 4 that $\iota_{1,3,1}({\texttt{L}^{\texttt{3}}})$ is parameterized by $f$ in Section 1.6. Thus, by Section 2, it suffices to show (3.5) for $n=1,2$ for $h=2,3$. This holds by Table 1 and the result follows. ∎
Remark 3.9 (Rational scrolls as curves in $\operatorname{Gr}(1,\mathbb{P}^{N})$).
The ideas used in the proof above can be put into a broader framework. In this remark, we provide an additional example to motivate the study of ED degrees of subvarieties of the Grassmannian.
A Bezier curve in $\mathbb{R}^{N}$ of degree $E$ is defined by a set of control points $P_{0},\dots,P_{E}\in\mathbb{K}^{N}$. The curve is given parametrically by
$\mathbf{B}(s,t)=\sum_{i=0}^{E}B_{i,E}(s,t)\cdot{P}_{i},\quad s=1,\,t\in[0,1]$
where $B_{i,E}(s,t)$ are the homogenized Bernstein polynomials $B_{i,E}(s,t)=\binom{E}{i}(s-t)^{E-i}t^{i}.$ This parameterization has the advantage that the curve segment $\{\mathbf{B}(1,t):t\in[0:1]\}$ on the standard affine chart in $\mathbb{R}^{N}$ is in the convex hull of the control points, starts at $P_{0}$, and ends at $P_{E}$. It follows from Theorem 2.3 that if the control points are in generic position, then $\operatorname{affEDdeg}(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\mathbf{B}(s,t))=En-2$ for a generic camera configuration $\mathbf{C}$.
Suppose now we have a pair of Bezier curves $\mathbf{B}_{i}(1,t)$, $i=1,2$, in $\mathbb{K}^{3}$ with degree $E_{i}$. We obtain a one-dimensional family of lines by considering the line spanned by $\mathbf{B}_{1}(1,t)$ and $\mathbf{B}_{2}(1,t)$ for $t\in[0,1]$. This is visualized in Figure 5. Again we assume the control points are in generic position. The Zariski closure of the family of lines is a subvariety of $\operatorname{Gr}(1,\mathbb{P}^{3})$, and we denote it as $\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}}$. The Plücker embedding of this variety is given by the two by two minors of the $2\times 4$ matrix
$\begin{bmatrix}s^{E_{1}}&\mathbf{B}_{1}(s,t)\\ s^{E_{2}}&\mathbf{B}_{2}(s,t)\end{bmatrix}$.
This gives a degree $E_{1}+E_{2}$ polynomial map from $\mathbb{P}^{1}$ to $\mathbb{P}^{5}$, and the results of Theorem 2.3 imply
$\operatorname{affEDdeg}(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}})\,=\,3(E_{1}+E_{2})n-2$
where $\mathbf{C}$ is an arrangement of $n$ cameras of size $3\times 4$.
Figure 5: The purple ruled surfaces consist of line segments connecting a pair of Bezier curves of degree one (left) and four (right). The lines spanned by the segments in a ruling give $\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}}\subset\operatorname{Gr}(1,\mathbb{P}^{3})$. The left surface has two rulings as discussed in Section 1.6.
4 Conclusion and future directions
In this article, we considered one-dimensional multiview varieties. We used topological arguments to give the first theoretical results for ED degrees of multiview varieties anchored at a Schubert variety. One of the key features of our proof is that each genericity condition is the complement of a union of hypersurfaces whose defining equations only rely on a pair of cameras or a single camera. This resulted in a surprising Section 2 to determine the ED degrees of multiview varieties where cameras have a prescribed structure. This is relevant for families of “dual cameras” found in resectioning [4] as well as calibrated cameras [14, 19] used in practice. We leave for future work the task of of generalizing these ideas to higher dimensional multiview varieties. This becomes more difficult because the camera centers may intersect the world variety. We also leave for future work the problem of determining the ED degree of line multiview varieties using a different embedding into the Grassmannian [16, 17, 5].

### v2

@@
-The Euclidean distance degree of curves: from rational to line multiview varieties
+The Euclidean distance degree of one-parameter anchored multiview varieties
 Abstract
-The Euclidean distance (ED) degree is an invariant that measures the algebraic complexity of optimizing the distance function of a point to a model.
-It has been studied in algebraic statistics, machine learning, and computer vision.
+Multiview varieties are mathematical models for the set of image feature correspondences that can be produced by a given camera arrangement.
+They possess an invariant known as their Euclidean distance (ED) degree, which measures the algebraic complexity of determining the 3D features that minimize the reprojection error when computing the scene structure by triangulation.
 In this article, we prove a formula for the ED degree of curves parameterized by rational functions with mild genericity assumptions.
@@
 Introduction
-One of the most impactful applications of algebraic geometry is in computer vision.
-This has led to the development of what is now known as algebraic vision.
-A featured problem in this area is to determine the number of critical points of the triangulation and, more generally, the multiview triangulation problem in multiple view geometry [13].
-This amounts to minimizing the Euclidean distance function to an algebraic set known as the (affine) multiview variety.
-Recent work by Duff and Rydell introduced an extensive catalogue of multiview varieties motivated by new vision applications.
+One of the most impactful applications of algebraic geometry is in computer vision, leading to the development of the field now known as algebraic vision.
+A featured problem in this area is to determine the number of critical points arising in triangulation and, more generally, in multiview triangulation in multiple view geometry (see [19] for a textbook reference).
+This amounts to minimizing a least squares reprojection error, and is mathematically framed as minimizing the squared Euclidean distance function to an algebraic set known as the (affine) multiview variety.
+Recent work by Duff and Rydell introduced an extensive catalogue of multiview varieties motivated by reconstruction applications [12, 18, 35, 31] and recent theoretical advances [4, 5, 20, 32, 33].
 In turn, this led to questions and conjectures on the Euclidean distance degrees for these models.
@@
 We also highlight prior work in the algebraic vision community to put our results into context and address some of the subtleties in the proofs.
-In Section 2 we prove a general statement on the Euclidean distance degree of the multiview variety of rational curves of arbitrary degree.
-In Section 3 we have the main application of our result: strengthening the catalogue theorem [7, Theorem 3.3] of Duff-Rydell by settling their conjectures on one-dimensional multiview varieties.
+In Section 2 we prove a general statement on the Euclidean distance degree of the multiview variety of a rational curve of arbitrary degree.
+In Section 3 we have the main application of our result: settling the conjectures [9, Conjecture 7.4.5 and Conjecture 7.4.6] of Duff–Rydell on one-dimensional multiview varieties.
+In Section 4 we apply Section 2 to determine the ED degree of a class of one-parameter families of lines in the Grassmannian.
 In Section 5 we conclude with a brief discussion on future directions.
@@
 1.1 Multiprojective varieties and multidegrees
-A subvariety of a product of projective spaces is known as a multiprojective variety. Let $Z\subset(\mathbb{P}^{h})^{n}$ be such a variety. Its defining ideal lies in the multigraded ring
+Solution sets of polynomial equations are algebraic varieties.
+When each polynomial is homogeneous, the solution set is invariant under rescaling, which gives rise to a projective variety.
+When the defining equations are homogeneous with respect to multiple independent rescalings, we obtain a subvariety of a product of projective spaces known as a multiprojective variety.
+In this subsection we review some basics of multiprojective varieties and their multidegrees, and in the next section we see how they arose in computer vision as multiview varieties.
+Let $Z\subset(\mathbb{P}^{h})^{n}$ be such a variety. Its defining ideal lies in the multigraded ring
 $\mathbb{K}[\overline{\mathbf{x}}_{1},\dots,\overline{\mathbf{x}}_{n}]\quad\text{ where }\quad\overline{\mathbf{x}}_{i}=(x_{i,0},\dots,x_{i,h})$
 and $\mathbb{K}$ is a field. If $Z$ is irreducible and of dimension $\dim{Z}$, then the multidegree of $Z$ is the polynomial
-$\operatorname{Deg}(Z):=\sum_{\alpha_{1}+\cdots+\alpha_{n}=\dim{Z}}c_{\alpha}\frac{T_{1}^{h}\cdots T_{n}^{h}}{T^{\alpha}}$
+$\operatorname{Deg}(Z):=\sum_{\alpha_{1}+\cdots+\alpha_{n}=\dim{Z}}c_{\alpha}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{T_{1}^{h-\alpha_{1}}\cdots T_{n}^{h-\alpha_{n}}}}$
 in the quotient ring $\mathbb{Z}[T_{1},\dots,T_{n}]/\langle T_{1}^{h+1},\dots,T_{n}^{h+1}\rangle$.
-We use the standard multi-index notation here, so that if $\alpha=(\alpha_{1},\dots,\alpha_{n})$ then $T^{\alpha}:=T_{1}^{\alpha_{1}}\cdots T_{n}^{\alpha_{n}}$.
+This quotient reflects the geometric fact that one cannot have a nonempty intersection of a factor $\mathbb{P}^{h}$ with more than $h$ general hyperplanes.
+We will use the standard multi-index notation, so that if $\alpha=(\alpha_{1},\dots,\alpha_{n})$ then $T^{\alpha}:=T_{1}^{\alpha_{1}}\cdots T_{n}^{\alpha_{n}}$.
 The coefficient $c_{\alpha}$ is the number of intersection points of $Z$ with a general linear space of the form
@@
 where $\ell_{i,j}$ is a general linear form in the unknowns $\overline{\mathbf{x}}_{i}$.
+One should think of the exponents of the terms of the polynomial as recording how many hyperplanes we impose in each factor, and of the coefficients as counting the size of the corresponding intersection.
 If $Z$ is reducible, then the multidegree of $Z$ is defined to be the sum of the multidegrees of the irreducible components.
@@
 Example 1.1.
-Consider the bi-homogeneous polynomials
+The following shows how multidegrees behave under intersecting bihomogeneous hypersurfaces. Consider the bi-homogeneous polynomials
 $f_{d}=x_{0}y_{0}^{d}+x_{1}y_{1}^{d}+x_{2}y_{2}^{d}+x_{3}y_{3}^{d},\quad d=0,1,2,3.$
@@
 $(T_{1}+1T_{2})(T_{1}+2T_{2})(T_{1}+3T_{2})=T_{1}^{3}+6T_{1}^{2}T_{2}+11T_{1}T_{2}^{2}+6T_{2}^{3}$
-and the multidegree of $V(f_{1},f_{2},f_{3})\cap V(f_{0})$ is $6T_{1}^{3}T_{2}+11T_{1}^{2}T_{2}^{2}+6T_{1}T_{2}^{3}.$ $\diamond$
+and the multidegree of $V(f_{1},f_{2},f_{3})\cap V(f_{0})$ is $6T_{1}^{3}T_{2}+11T_{1}^{2}T_{2}^{2}+6T_{1}T_{2}^{3}.$
+This example shows how multidegrees behave under intersecting bihomogeneous hypersurfaces. $\diamond$
 1.2 Multiprojective versus affine multiview varieties
@@
 However, optimization problems such as Euclidean distance minimization are typically formulated in an affine chart.
-We now describe both perspectives to get a multiprojective and an affine multiview variety.
+We now describe how multiview varieties appear in both multiprojective and affine settings.
+In computer vision, these encode the algebraic constraints satisfied by feature correspondences across multiple images.
 An arrangement $\mathbf{C}=(C_{1},\dots,C_{n})$ of full rank $3\times 4$ matrices, called cameras, gives a multiprojective multiview variety of $\mathbb{P}^{3}$ as the Zariski closure of the image of the rational map
@@
 $z$ $\mapsto(C_{1}z,\dots,C_{n}z).$
-As in Section 1.1, we take the coordinates of $(\mathbb{P}^{2})^{n}$ to be $\overline{\mathbf{x}}_{1},\dots,\overline{\mathbf{x}}_{n}$ The multidegree of the image of (1.1) is
-$\sum_{\max\{\alpha_{1},\dots,\alpha_{n}\}=2,\,\,\alpha_{1}+\cdots+\alpha_{n}=3}1\cdot{T_{1}^{2-\alpha_{1}}\cdot T_{2}^{2-\alpha_{2}}\cdot\cdots\cdot T_{n}^{2-\alpha_{n}}},\quad n>1.$ (1.2)
-For a textbook introduction to multidegrees see [23, Chapter 8]. Ideals of multiview varieties are also well studied in algebraic vision, and for more details we highlight [1, 3, 9].
+As in Section 1.1, we take the coordinates of $(\mathbb{P}^{2})^{n}$ to be $\overline{\mathbf{x}}_{1},\dots,\overline{\mathbf{x}}_{n}$. The multidegree of the image of (1.1) is
+$\sum_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\max\{\alpha_{1},\dots,\alpha_{n}\}\leq 2,}\,\,\alpha_{1}+\cdots+\alpha_{n}=3}1\cdot{T_{1}^{2-\alpha_{1}}\cdot T_{2}^{2-\alpha_{2}}\cdot\cdots\cdot T_{n}^{2-\alpha_{n}}},\quad n>1.$ (1.2)
+The exponent $2-\alpha_{i}$ indicates that $\alpha_{i}$ hyperplanes were used in the $i$th factor.
+Ideals and multidegrees of multiview varieties are also well studied in algebraic vision, and for more details we highlight [1, 5, 11].
+For a textbook introduction to multidegrees, see [30, Chapter 8].
 The affine multiview variety of $\mathbb{P}^{3}$ is obtained by restricting each factor of $(\mathbb{P}^{2})^{n}$ to an affine chart.
@@
 $z\mapsto\begin{bmatrix}\displaystyle{\frac{C_{1}^{(1)}\cdot z}{C_{1}^{(0)}\cdot z}}&\dots&\displaystyle{\frac{C_{n}^{(1)}\cdot z}{C_{n}^{(0)}\cdot z}}\\ \\ \displaystyle{\frac{C_{1}^{(2)}\cdot z}{C_{1}^{(0)}\cdot z}}&\dots&\displaystyle{\frac{C_{n}^{(2)}\cdot z}{C_{n}^{(0)}\cdot z}}\end{bmatrix}\in\mathbb{K}^{2\times n}.$
-In general, if $Z$ is a multiprojective variety then we denote its restriction to the conventional affine chart by $Z_{{\operatorname{Aff}}}\subset(\mathbb{K}^{h})^{n}$.
+In general, if $Z$ is a multiprojective subvariety in $(\mathbb{P}^{h})^{n}$ then we denote its restriction to the conventional affine chart by $Z_{{\operatorname{Aff}}}\subset(\mathbb{K}^{h})^{n}$.
+This corresponds to working with inhomogeneous image coordinates, where reprojection error is naturally defined.
 1.3 Euclidean distance degree
 The Euclidean distance degree (ED degree) of an algebraic variety measures the algebraic complexity of the problem of minimizing Euclidean distance to that variety.
-This invariant plays a central role in applications such as computer vision, statistics, and optimization.
 We recall its definition in this subsection.
@@
 Namely, for a smooth point $x\in X_{\mathbb{R}}$, we require $(x-u)^{T}v=0$ for all $v$ in the tangent space $T_{x}X_{\mathbb{R}}$ of $X_{\mathbb{R}}$ at $x$.
-This condition gives rise to a polynomial system [6, Equation 2.1] that has $x$ as the unknown and $u$ as the given generic data.
+This condition gives rise to a polynomial system [8, Equation 2.1] that has $x$ as the unknown and $u$ as the given generic data.
 The number of complex (real or nonreal) solutions to this polynomial system is called the Euclidean distance degree of $X_{\mathbb{R}}$ [8].
-The ED degree is a measure of complexity for minimizing the Euclidean distance from $X_{\mathbb{R}}$ to $u$.
+The ED degree is a measure of complexity for minimizing the squared Euclidean distance from $X_{\mathbb{R}}$ to $u$.
 Example 1.2.
@@
 $g_{1}=x^{2}+y^{2}+z^{2}-1,\quad g_{2}=x^{4}+y^{4}+z^{4}+ax^{2}y^{2}+by^{2}z^{2}+cz^{2}x^{2}-s\left(x^{2}+y^{2}+z^{2}\right)^{2}$
-where (a,b,c,s)=(0.8,0.74,0.5,0.32)(a,b,c,s)=(-0.8,-0.74,-0.5,0.32). Then, for a generic data point $u$, the ideal of the critical points is generated by $g_{1},g_{2}$ and the determinant of
+where (a,b,c,s)=(0.8,0.74,0.5,0.32)(a,b,c,s)=(-0.8,-0.74,-0.5,0.32).
+Then, for a generic data point $u$, the ideal of the critical points is generated by $g_{1},g_{2}$ and the determinant of the $3\times 3$ matrix
 $\begin{bmatrix}\nabla d_{u}&\nabla g_{1}&\nabla g_{2}\end{bmatrix}$
@@
 Note, when $X$ is singular an additional step of saturation needs to be performed to obtain the ideal of critical points. $\diamond$
-Figure 2: The red curve is $X_{\mathbb{R}}$ and the real ED critical points for the distance function to the orange points is the set of intersection points with the silver curve.
+Figure 2: The red curve is $X_{\mathbb{R}}$ and the real ED critical points for the distance function to the orange point is the set of intersection points with the silver curve.
 Beyond its algebraic definition, the Euclidean distance degree also has a topological interpretation. For more details, we refer to the survey article [29].
@@
 $U_{\boldsymbol{\beta}}:=\mathbb{C}^{h}\setminus\{(z_{1},\dots,z_{h})\in\mathbb{C}^{h}:\sum_{i=1}^{h}(z_{i}-\beta_{i})^{2}+\beta_{0}=0\}.$
-For many of our purposes, $X$ is smooth and it suffices to consider the Euler-Poincare characteristic $\chi$.
+For many of our purposes, $X$ is smooth and it suffices to consider the Euler–Poincaré characteristic $\chi$.
 Theorem 1.4 ([27]).
@@
 where $U_{\boldsymbol{\beta}}:=\mathbb{C}^{h}\setminus\{(z_{1},\dots,z_{h})\in\mathbb{C}^{h}:\sum_{i=1}^{h}(z_{i}-\beta_{i})^{2}+\beta_{0}=0\}$.
-In Section 2 we will apply these results to compute the ED degree of one dimensional multiview varieties.
+We apply these results to compute the ED degrees of one-dimensional multiview varieties in Section 2 and resolve computer vision conjectures in Section 3.
 1.4 The projective ED degree
@@
 In this subsection, we recall this invariant and prior work for curves.
+Although our main results concern affine ED degrees, the projective ED degree also plays an important role because many multiview varieties arise as affine cones or have natural projective closures.
 Let $Z\subset\mathbb{P}^{h}$ be an irreducible closed subvariety of codimension $c$ and let $I_{Z}$ be its defining ideal.
-If the homogeneous polynomials $g_{1},\dots,g_{s}\in\mathbb{C}[z_{0},z_{1},\dots,z_{h}]$ generate $I_{Z}$, then the projective Euclidean distance degree correspondence is a biprojective variety in $\mathbb{P}^{h}\times\mathbb{P}^{h}$ defined by the saturated ideal
+If the homogeneous polynomials $g_{1},\dots,g_{s}\in\mathbb{C}[z_{0},z_{1},\dots,z_{h}]$ generate $I_{Z}$, then the projective Euclidean distance correspondence is a biprojective variety in $\mathbb{P}^{h}\times\mathbb{P}^{h}$ defined by the saturated ideal
 $\biggl(I_{Z}+\biggl\langle\hbox{$(c+2)\times(c+2)$-minors of $\begin{pmatrix}u\\ z\\ J(g)\end{pmatrix}$}\biggr\rangle\biggr):\bigl(I_{Z_{\rm sing}}\cdot\langle z_{0}^{2}+\cdots+z_{h}^{2}\rangle\,\bigr)^{\infty}$ (1.3)
@@
 $I_{Z_{\operatorname{sing}}}\,\,=\,\,I_{Z}+\bigl\langle\hbox{$(c\times c)$-minors of $J(g)$}\bigr\rangle.$
-The ED bidegree is defined to be multidegree of the ED degree correspondence.
-The coefficient of $T_{2}^{h}$ is the (projective) ED degree of $X$, and is denoted by $\operatorname{EDdeg_{proj}}(X)$.
-The other coefficients of the multidegree are known as the sectional ED degree, and the degree of $X$ is the coefficient of $T_{1}^{c}T_{2}^{h-c}$.
+The ED bidegree is defined to be multidegree of the ED correspondence. The coefficient of $T_{2}^{h}$ is the (projective) ED degree of $X$, and is denoted by $\operatorname{EDdeg_{proj}}(X)$.
 The affine cone over $Z$ is
@@
 It also allows the use of several topological formulas.
-Theorem 1.5 (Theorem 1.3 [21]).
-Let $Z$ be any irreducible closed subvariety of $\mathbb{P}^{h}$. Then for a general ${{\boldsymbol{\beta}}}=(\beta_{0},\ldots,\beta_{h})\in{{\mathbb{C}}}^{{h+1}}\setminus\{0\}$ we have
+Theorem 1.5.
+[28, Theorem 1.3] Let $Z$ be any irreducible closed subvariety of $\mathbb{P}^{h}$.
+Then for a general ${{\boldsymbol{\beta}}}=(\beta_{0},\ldots,\beta_{h})\in{{\mathbb{C}}}^{{h+1}}\setminus\{0\}$ we have
 $\operatorname{EDdeg_{proj}}(Z)=(-1)^{\dim X}\chi\big({\rm Eu}_{Z}|_{\mathcal{U}_{{\boldsymbol{\beta}}}}\big),$
@@
 Example 1.7 (Standard grading giving discrepancy).
-Consider the curve $Z$ in $\mathbb{P}^{2}\times\mathbb{P}^{2}$ parameterized by
-$\mathbb{P}^{1}\to Z\subset\mathbb{P}^{2}\times\mathbb{P}^{2}$
+Consider the curve $Z$ in $\mathbb{P}^{2}\times\mathbb{P}^{2}$ parameterized by the map
+$\mathbb{P}^{1}\to Z\subset\mathbb{P}^{2}\times\mathbb{P}^{2},$
 […]
@@
 $\diamond$
+The ED degree is not invariant under projective closure. Even for classical multiview varieties, the affine and projective versions may have drastically different ED degrees.
 Example 1.8 (Projective closure discrepancy).
 The affine $n$-view variety discussed in Section 1.2 is known [27] to have ED degree
-$\frac{9}{2}n^{3}-\frac{21}{2}n^{2}+8n-4.$
-However, by [12, Theorem 1], its projective closure in $\mathbb{P}^{2n}$ has ED degree
+$\frac{9}{2}n^{3}-\frac{21}{2}n^{2}+8n-4.$ (1.5)
+This cubic polynomial in $n$ counts the number of (real or nonreal) critical points of the distance function from a general data point to an $n$-view variety when the cameras are in generic position.
+The values of this polynomial for small $n$ was determined by [34].
+In particular, for $n=3$, there are forty seven critical points, and [34] showed how to minimize the distance function to a three-view variety by solving a $47\times 47$ eigenvalue problem.
+In [8], it was noticed that the projective closure in $\mathbb{P}^{5}$ of a three-view variety has ED degree 56.
+In [17, Theorem 1], they found the projective closure of the $n$-view variety in $\mathbb{P}^{2n}$ has ED degree
 $6n^{3}-15n^{2}+11n-4.\vskip-24.0pt$
@@
 1.5 More multiview varieties and triangulation
-In this subsection, we define what are known as anchored point multiview varieties. To stay organized and minimize notation, we focus on the camera arrangement and the image it induces.
+Anchored multiview varieties arise naturally when the world contains known structures, such as a curve, line, or surface.
+These models describe all possible image observations of points constrained to lie on $Y$.
+In this subsection, we recall the definition of anchored point multiview varieties.
+To stay organized and minimize notation, we focus on the camera arrangement and the image it induces.
 Definition 1.9.
@@
 $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\mathbb{P}^{N}:=\overline{\operatorname{im}(\Phi_{\mathbf{C}})}\subset(\mathbb{P}^{h})^{n},$
-where the overline denotes the Zariski closure. If $Y\subseteq\mathbb{P}^{N}$ is a subvariety, we define
+where the overline denotes the Zariski closure. If $Y\subseteq\mathbb{P}^{N}$ is a subvariety, define
 $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y:=\overline{\Phi_{\mathbf{C}}(Y)}$
@@
 When $N=3$ and $h=2$, the map $\Phi_{\mathbf{C}}$ models the procedure of taking $n$ pictures with $n$ possibly distinct cameras.
-A point $P\in\mathbb{P}^{3}$ is mapped to the tuple of images $(C_{1}P,\dots,C_{n}P)$ provided that $P$ is not in the kernel of any $C_{i}$.
+A world point $P\in\mathbb{P}^{3}$ is mapped to the tuple of images $(C_{1}P,\dots,C_{n}P)$ provided that $P$ is not in the kernel of any $C_{i}$.
 Points $P\in\mathbb{P}^{N}$ that lie in the kernel of some camera are referred to as being at a camera center, where $\Phi_{\mathbf{C}}$ is undefined.
 That is, the camera centers constitute the base locus of $\Phi_{\mathbf{C}}$.
-Example 1.10 (A one dimensional multiview variety).
+Example 1.10 (A one-dimensional multiview variety).
 If $Y$ is a line in $\mathbb{P}^{3}$, consider the anchored multiview variety $(C_{1},\dots,C_{n})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y$.
 The affine multiview variety of $(C_{1},\dots,C_{n})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y$ has ED degree 3n23n-2.
-This result appears in [24, Theorem 1.7] and serves as a baseline case for understanding the complexity of triangulating points on curves.
+This result appears in [33, Theorem 1.7] and serves as a baseline case for understanding the complexity of triangulating points on curves.
+This is the degree one case of Theorem 2.3.
 $\diamond$
@@
 $\operatorname{affEDdeg}(X):=\operatorname{EDdeg}(X_{\operatorname{Aff}})$ (1.7)
-where $X_{\operatorname{Aff}}$ is the restriction of $X$ to the conventional affine chart described in Section 1.2.
+where $X_{\operatorname{Aff}}$ is the restriction of $X$ to the affine chart described in Section 1.2.
 1.6 A Schubert variety in $\operatorname{Gr}(1,\mathbb{P}^{3})$
-In this subsection, we define a subvariety of the Grassmannian that will induce the multiprojective varieties appearing in our results in Section 3.
+In this subsection, we consider a subvariety of the Grassmannian that will induce the multiprojective varieties appearing in our results in Section 3.
 Let $\operatorname{Gr}(1,\mathbb{P}^{3})$ denote the Grassmannian of lines in $\mathbb{P}^{3}$.
-Given three skew lines ${\ell^{(1)}},{\ell^{(2)}},{\ell^{(3)}}$ we define
+Given three skew lines ${\ell^{(1)}},{\ell^{(2)}},{\ell^{(3)}}$ we have
 ${\texttt{L}^{\texttt{3}}}({\ell^{(1)}},{\ell^{(2)}},{\ell^{(3)}}):=\{{\ell}\in\operatorname{Gr}(1,\mathbb{P}^{3}):{\ell}\cap{\ell^{(i)}}\neq\emptyset,\,i=1,2,3\}.$ (1.8)
@@
 The Plücker embedding
-$\iota:\operatorname{Gr}(1,\mathbb{P}^{3})\to\mathbb{P}^{5},$
+$\iota:\operatorname{Gr}(1,\mathbb{P}^{3})\to\mathbb{P}^{5}$
 provides us a way to embed ${\texttt{L}^{\texttt{3}}}\subset\operatorname{Gr}(1,\mathbb{P}^{3})$ into $\mathbb{P}^{5}$.
@@
 $\nu([s:t])=\begin{bmatrix}s^{2}\\ st\\ t^{2}\end{bmatrix}$
-with a linear map $B_{\texttt{L}^{\texttt{3}}}$ represented by the matrix
-$B_{{\texttt{L}^{\texttt{3}}}}=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&-1&0\\ 0&1&0\\ 0&0&0\\ 0&0&1\end{bmatrix}$
-so that the parameterization map $f:\mathbb{P}^{1}\to\iota({\texttt{L}^{\texttt{3}}})$ takes $[s:t]$ to $B_{{\texttt{L}^{\texttt{3}}}}\cdot\nu([s:t])$.
+with a linear map, denoted by $B_{\texttt{L}^{\texttt{3}}}$, so that the parameterization map $f:\mathbb{P}^{1}\to\iota({\texttt{L}^{\texttt{3}}})$ takes $[s:t]$ to $B_{{\texttt{L}^{\texttt{3}}}}\cdot\nu([s:t])$.
 Note, our choice for the three skew lines ${\ell^{(1)}},{\ell^{(2)}},{\ell^{(3)}}$ in (1.9) was made without loss of generality as the unique quadric containing any other choice of three skew lines is projectively equivalent to $V(x_{0}x_{3}-x_{1}x_{2})$.
@@
 Remark 1.14.
-If $\mathbf{D}$ is a camera arrangement of size $(H+1)\times 6$, then $\mathbf{D}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\iota({\texttt{L}^{\texttt{3}}})\subset(\mathbb{P}^{H})^{n}$ is a point multiview variety anchored at $\iota({\texttt{L}^{\texttt{3}}})$.
+If $\mathbf{D}$ is an arrangement of cameras of size $(H+1)\times 6$, then $\mathbf{D}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\iota({\texttt{L}^{\texttt{3}}})\subset(\mathbb{P}^{H})^{n}$ is a point multiview variety anchored at $\iota({\texttt{L}^{\texttt{3}}})$.
 This variety was previously studied in [9] when $\mathbf{D}$ is an arrangement of so-called wedge cameras (Section 3.2).
 We revisit this construction and show how it relates to line multiview varieties in Section 3 for $H=\binom{2+1}{2}-1$ and $H=\binom{3+1}{2}-1$.
-2 ED degree of multiview varieties of a rational curve
+2 ED degrees of multiview varieties anchored at rational curves
 In this section, we let $Y$ be a degree $E$ rational curve in $\mathbb{P}^{N}$ with $N\geq 3$ parametrized by
@@
 Proof.
-For $n=1$, this is a standard algebraic geometry fact. For instance, see [8, Exercise 3.34]. The case when $n>1$ immediately follows. ∎
+For $n=1$, this is a standard algebraic geometry fact. For instance, see [10, Exercise 3.34]. The case when $n>1$ immediately follows. ∎
 Example 2.2 (Single view varieties of cubic curves).
@@
 $V(C_{i}^{(0)}\cdot f,C_{j}^{(0)}\cdot f)=\emptyset$ (2.5)
-for all $i,j\in[n],\ i\neq j$, then $S_{\infty}$ has $ne$ points.
-This holds by genericity of the camera arrangement: the degree $e$ polynomial $C_{i}^{(0)}\cdot f$ has a multiple root if and only if its $X$-discriminant vanishes, and $C_{i}^{(0)}\cdot f,C_{j}^{(0)}\cdot f$ share a root if and only if their resultant is nonvanishing [11, Section 1.1].
-Therefore, if $f$ is a generically one-to-one map then $H_{\infty}\cap X$ has precisely $ne$ points.
+for all $i,j\in[n],\ i\neq j$, then $S_{\infty}$ has $En$ points.
+This holds by genericity of the camera arrangement: the degree $E$ polynomial $C_{i}^{(0)}\cdot f$ has a multiple root if and only if its $X$-discriminant vanishes, and $C_{i}^{(0)}\cdot f,C_{j}^{(0)}\cdot f$ share a root if and only if their resultant is nonvanishing [14, Section 1.1].
+Therefore, if $f$ is a generically one-to-one map then $H_{\infty}\cap X$ has precisely $En$ points.
 For $\beta=(\beta_{0},\beta_{1,1},\dots,\beta_{n,h})\in\mathbb{R}^{nh+1}$, define the hypersurface
@@
 $g_{\beta}=\beta_{0}D+\sum_{i=1}^{n}\frac{D}{\left(C_{i}^{(0)}\cdot f\right)^{2}}\sum_{j=1}^{h}\left({(C_{i}^{(1)}\cdot f)}-\beta_{i,j}{(C_{i}^{(0)}\cdot f})\right)^{2}$
-is a homogeneous degree $2ne$ polynomial in $s,t$.
-Because $\beta_{0}$ is generic, if $g_{\beta}$ has no common roots with $D$ then $g_{\beta}$ has $2ne$ distinct roots.
+is a homogeneous degree $2En$ polynomial in $s,t$.
+Because $\beta_{0}$ is generic, if $g_{\beta}$ has no common roots with $D$ then $g_{\beta}$ has $2En$ distinct roots.
 On the other hand, if $g_{\beta}$ had a common root with $D$, then there would exist a unique $i=a$ such that $\left(C_{i}^{(0)}\cdot f\right)$ vanishes on that root.
@@
 Thus, in an $n$-camera configuration, $\chi(X_{\operatorname{Aff}}\cap Q_{\beta})=2En$.
-If $X_{\operatorname{Aff}}$ is smooth, an application of [20, Theorem 3.8] implies that
+If $X_{\operatorname{Aff}}$ is smooth, an application of [27, Theorem 3.8] implies that
 $\operatorname{affEDdeg}(X)=-(\chi(X)-\chi(X\cap H_{\infty})-\chi(X\cap Q_{\beta})).$ (2.8)
@@
 Suppose $f$ is as in (2.1) with $N\geq 3$ and $h\geq 2$.
-Let $\mathcal{C}\subset\mathbb{P}^{(h+1)\times(N+1)-1}$ be an irreducible projective variety.
-If $C_{1},\dots,C_{n}$ are generic points in $\mathcal{C}$ and
+Let $\mathcal{A}\subset\mathbb{P}^{(h+1)\times(N+1)-1}$ be an irreducible projective variety.
+If $C_{1},\dots,C_{n}$ are generic points in $\mathcal{A}$ and
 $\operatorname{affEDdeg}((C_{1},\dots,C_{n})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1}))=3En-2$ (2.11)
@@
 Define the following irreducible variety of cameras:
-$\mathcal{C}:=\left\{\begin{bmatrix}c_{1}&c_{2}&0&0&0&0\\ 0&0&c_{3}&c_{4}&0&0\\ 0&0&0&0&c_{5}&c_{6}\end{bmatrix}\in\mathbb{K}^{3\times 6}:c_{i}\in\mathbb{K}\right\}.$
-If $C$ is a generic element in $\mathcal{C}$, then $\operatorname{affEDdeg}(C_{1}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y)=9$. We see that this disagrees with the numbers in (2.11).
+$\mathcal{A}:=\left\{\begin{bmatrix}c_{1}&c_{2}&0&0&0&0\\ 0&0&c_{3}&c_{4}&0&0\\ 0&0&0&0&c_{5}&c_{6}\end{bmatrix}\in\mathbb{K}^{3\times 6}:c_{i}\in\mathbb{K}\right\}.$
+If $C$ is a generic element in $\mathcal{A}$, then $\operatorname{affEDdeg}(C_{1}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y)=9$. We see that this disagrees with the numbers in (2.11).
 $\diamond$
@@
 Consider the following family of $3\times 4$ camera matrices:
-$\mathcal{C}=\left\{[c_{i,j}]\in\mathbb{K}^{3\times 4}:c_{1,1}=0\right\}.$
-Let $C_{1}$ and $C_{2}$ be generic points of $\mathcal{C}$. If $Y$ is the rational normal curve in $\mathbb{P}^{3}$ parameterized by $t^{3}:st^{2}:s^{2}t:s^{3}]$, then
+$\mathcal{A}=\left\{[c_{i,j}]\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{P}^{3\times 4-1}}:c_{1,1}=0\right\}.$
+Let $C_{1}$ and $C_{2}$ be generic points of $\mathcal{A}$. If $Y$ is the rational normal curve in $\mathbb{P}^{3}$ parameterized by $t^{3}:st^{2}:s^{2}t:s^{3}]$, then
 $\operatorname{affEDdeg}(C_{1}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y)=7\text{ and }\operatorname{affEDdeg}((C_{1},C_{2})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y)=13.$
@@
 The projective ED degrees of $X$ and $C\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}X$ both equal $3\cdot 4-2$.
-As affine cones in ${{\mathbb{C}}}^{4+1}$ and ${{\mathbb{C}}}^{3}$, these varieties are parameterized by degree $4$ polynomials and one might predict from [6, Section 2] that the ED degree is $(2\cdot 4-1)^{2}$.
+As affine cones in ${{\mathbb{C}}}^{4+1}$ and ${{\mathbb{C}}}^{3}$, these varieties are parameterized by degree $4$ polynomials and one might predict from [8, Section 2] that the ED degree is $(2\cdot 4-1)^{2}$.
 The reason we do not get this bound is because these are general homogeneous polynomials of degree $4$ and not general degree $4$ polynomials.
@@
 Remark 2.9 (Intersection theory and multidegrees).
-The results in this section have an intersection theoretic [10] interpretation through the multigrees presented in Section 1.2.
-Namely, the proof in Theorem 2.3 is showing that the curve $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})\subset(\mathbb{P}^{h})^{n}$ has multidegree $\sum_{i=1}^{n}e\cdot\frac{T_{1}^{h}T_{2}^{h}\cdots T_{n}^{h}}{T_{i}}$ and intersects transversally each of the hyperplanes $V(x_{i,0})$ for $i=1,\dots,n$.
+The results in this section have an intersection theoretic [13] interpretation through the multidegrees presented in Section 1.2.
+Namely, the proof in Theorem 2.3 is showing that the curve $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})\subset(\mathbb{P}^{h})^{n}$ has multidegree $\sum_{i=1}^{n}E\cdot{T_{1}^{h-1}T_{2}^{h-1}\cdots T_{n}^{h-1}}$ and intersects transversally each of the hyperplanes $V(x_{i,0})$ for $i=1,\dots,n$.
 This implies $\chi(X\cap H_{\infty})=En$.
@@
 (obtained by homogenizing the polynomial in (2.6)) also transversally intersects the curve $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})$ with no solutions at infinity.
-Since the multidegree of $V(q_{\beta})$ is $\sum_{i=1}^{n}2T_{i}^{2}$, we have $\chi(X\cap V(q_{\beta}))=2en$ because
-$\left(\sum_{i=1}^{n}2T_{i}^{2}\right)\cdot\left(\sum_{i=1}^{n}e\cdot\frac{T_{1}^{h}\cdots T_{n}^{h}}{T_{i}}\right)=2en\cdot T_{1}^{h}\cdots T_{n}^{h}\in\mathbb{Z}[T_{1},\dots,T_{n}]/\langle T_{1}^{h+1},\dots,T_{n}^{h+1}\rangle.$
-3 Strengthening the ED-equivalence catalogue
-In this section, we present our main application of our results.
+Since the multidegree of $V(q_{\beta})$ is $\sum_{i=1}^{n}2T_{i}$, we have $\chi(X\cap V(q_{\beta}))=2En$ because
+$\left(\sum_{i=1}^{n}2T_{i}\right)\cdot\left(\sum_{i=1}^{n}E\cdot{T_{1}^{h-1}\cdots T_{n}^{h-1}}\right)=2En\cdot T_{1}^{h}\cdots T_{n}^{h}\in\mathbb{Z}[T_{1},\dots,T_{n}]/\langle T_{1}^{h+1},\dots,T_{n}^{h+1}\rangle.$
+3 Settling conjectures of Duff–Rydell
+In this section, we present our main application of our Theorem 2.3.
 Section 3.1 defines our main varieties of interest: a set of multiview varieties anchored at a subvariety of the Grassmannian.
@@
 Section 1.5 considers the point multiview variety $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y$ anchored at $Y$ as the Zariski closure of the image of the map (1.6) for $Y\subset\mathbb{P}^{N}$.
-In this section, we carry this idea over to a linear map between Grassmannians.
+In this section, we carry this idea over to a map between Grassmannians.
 The associated multiview variety (called a generalized multiview variety in [9]) is obtained as the Zariski closure of a rational map
@@
 In [9], the full data of a multiview variety is recorded in the notation $\mathcal{M}_{\mathbf{C},k}^{\Lambda,N,h}$.
-We opt for the more parsimonious notation and refer to the catalogue in [7, Theorem 3.3] for a complete description of multiview varieties arising from projections of points and lines in 1, 2, and 3-dimensional projective space.
+We opt for the more parsimonious notation and refer to the catalogue in [9, Theorem 3.3] for a complete description of multiview varieties arising from projections of points and lines in 1, 2, and 3-dimensional projective space.
 As defined, $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\Lambda$ is a subvariety of a product of Grassmannians.
@@
 h=2,n=1h=2,\,n=1 $\mathbb{P}^{2}$ $2T_{1}$ 4
-h=2,n=2h=2,\,n=2 $\mathbb{P}^{2}\times\mathbb{P}^{2}$ $2T_{1}+2T_{2}$ 10
-h=3,n=1h=3,\,n=1 $\mathbb{P}^{5}$ $2T_{1}$ 4
-h=3,n=2h=3,\,n=2 $\mathbb{P}^{5}\times\mathbb{P}^{5}$ $2T_{1}+2T_{2}$ 10
+h=2,n=2h=2,\,n=2 $\mathbb{P}^{2}\times\mathbb{P}^{2}$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2T_{1}+2T_{2})T_{1}T_{2}$} 10
+h=3,n=1h=3,\,n=1 $\mathbb{P}^{5}$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2T_{1}^{4}$} 4
+h=3,n=2h=3,\,n=2 $\mathbb{P}^{5}\times\mathbb{P}^{5}$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2T_{1}+2T_{2})T_{1}^{4}T_{2}^{4}$} 10
 Table 1: Summary of $\mathbf{C}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}{\texttt{L}^{\texttt{3}}}$ with $\mathbf{C}^{h}$ having $n=1,2$ cameras in its arrangement.
@@
 Next, we show how to realize $\mathbf{C}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}{\texttt{L}^{\texttt{3}}}$ as a point multiview variety anchored at $\iota({\texttt{L}^{\texttt{3}}})\subset\mathbb{P}^{5}$.
-To that end, we review the exterior algebra (see [18, Chapter XVI] for a textbook reference) to state Section 3.2.
+To that end, we review the exterior algebra (see [25, Chapter XVI] for a textbook reference) to state Section 3.2.
 Recall the exterior algebra of a vector space $V$ over a field $\mathbb{K}$ is the graded algebra
 $\bigwedge V:=\bigoplus_{k=0}^{\dim V}\bigwedge^{k}V,$
-where $\bigwedge^{k}V$ is the space of alternating $k$-linear forms on $V$.
-Elements of $\bigwedge^{k}V$ are called $k$-vectors, and the wedge product satisfies $u\wedge v=-v\wedge u$ and $v\wedge v=0$.
+where $\bigwedge^{k}V$ is the space of $k$-vectors, the dual of the space of alternating $k$-linear forms on $V$.
+The wedge product satisfies $u\wedge v=-v\wedge u$ and $v\wedge v=0$.
 A simple wedge is one of the form $v_{1}\wedge\cdots\wedge v_{k}$, $v_{i}\in V.$ Every element of $\bigwedge^{k}V$ is a linear combination of simple wedges.
@@
 $((1,2),(1,3),(2,3),(1,4),(2,4),(3,4))$ for columns. Then,
-${\wedge^{2}C}=\begin{bmatrix}-1&0&0&3&4&0\\ 0&5&0&10&0&-20\\ 0&0&0&-5&0&0\end{bmatrix}$.
+${\wedge^{2}C}=\begin{bmatrix}-1&0&3&0&4&0\\ 0&5&10&0&0&-20\\ 0&0&-5&0&0&0\end{bmatrix}$.
 Each entry is $\det C_{I,J}$ with $I$ a row pair and $J$ a column pair in the stated order. $\diamond$
@@
 Proof.
-Let $\mathbf{D}^{h}=(D_{1},\dots,D_{n})$ denote the $2$-wedge camera configuration of $\mathbf{C}^{h}$.
+Let $h=2$ or $h=3$, and let $\mathbf{D}^{h}=(D_{1},\dots,D_{n})$ denote the $2$-wedge camera configuration of $\mathbf{C}^{h}$.
 By Section 3.2 and Section 3.3, we have
@@
 It follows from Figure 4 that $\iota_{1,3,1}({\texttt{L}^{\texttt{3}}})$ is parameterized by $f$ in Section 1.6.
-Thus, by Section 2, it suffices to show (3.5) for $n=1,2$ for $h=2,3$.
+Thus, by Section 2, it suffices to show (3.5) for $n=1,2$.
 This holds by Table 1 and the result follows.
 ∎
-Remark 3.9 (Rational scrolls as curves in $\operatorname{Gr}(1,\mathbb{P}^{N})$).
-The ideas used in the proof above can be put into a broader framework. In this remark, we provide an additional example to motivate the study of ED degrees of subvarieties of the Grassmannian.
-A Bezier curve in $\mathbb{R}^{N}$ of degree $E$ is defined by a set of control points $P_{0},\dots,P_{E}\in\mathbb{K}^{N}$. The curve is given parametrically by
-$\mathbf{B}(s,t)=\sum_{i=0}^{E}B_{i,E}(s,t)\cdot{P}_{i},\quad s=1,\,t\in[0,1]$
-where $B_{i,E}(s,t)$ are the homogenized Bernstein polynomials $B_{i,E}(s,t)=\binom{E}{i}(s-t)^{E-i}t^{i}.$ This parameterization has the advantage that the curve segment $\{\mathbf{B}(1,t):t\in[0:1]\}$ on the standard affine chart in $\mathbb{R}^{N}$ is in the convex hull of the control points, starts at $P_{0}$, and ends at $P_{E}$.
-It follows from Theorem 2.3 that if the control points are in generic position, then $\operatorname{affEDdeg}(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\mathbf{B}(s,t))=En-2$ for a generic camera configuration $\mathbf{C}$.
-Suppose now we have a pair of Bezier curves $\mathbf{B}_{i}(1,t)$, $i=1,2$, in $\mathbb{K}^{3}$ with degree $E_{i}$.
-We obtain a one-dimensional family of lines by considering the line spanned by $\mathbf{B}_{1}(1,t)$ and $\mathbf{B}_{2}(1,t)$ for $t\in[0,1]$.
+4 One-parameter families of 3D lines
+We study an additional model to further motivate the investigation of ED Degrees of subvarieties of the Grassmannian.
+One important example comes from rational scrolls.
+These are surfaces that represent one-parameter families of 3D lines and arise naturally in computer vision applications.
+In this section, we consider an instance of such a family: lines sweeping along Bézier curves to provide a one-parameter line-view variety.
+This is obtained by mapping a subvariety of $\operatorname{Gr}(1,\mathbb{P}^{3})$ to $\operatorname{Gr}(1,\mathbb{P}^{2})^{n}$.
+For background (e.g., [15, Chapter 5]), recall a Bézier curve in $\mathbb{K}^{N}$ of degree $E$ is defined by a set of $E+1$ control points
+$P_{0},\dots,P_{E}\in\mathbb{K}^{N}$
+with $N\geq 2$. The curve is given parametrically by
+$\mathbf{B}(1,t)=\sum_{i=0}^{E}B_{i,E}(1,t)\cdot{P}_{i},\quad\,t\in[0,1]$ (4.1)
+where $B_{i,E}(s,t)$ is the $i$th homogenized Bernstein polynomial of degree $E$,
+$B_{i,E}(s,t):=\binom{E}{i}(s-t)^{E-i}t^{i}.$
+This parameterization has the advantage that the curve segment $\{\mathbf{B}(1,t):t\in[0,1]\}$ on the standard affine chart in $\mathbb{K}^{N}$ is in the convex hull of the control points, starts at $P_{0}=\mathbf{B}(1,0)$, and ends at $P_{E}=\mathbf{B}(1,1)$.
+Using the homogenized polynomials, the parameterization from (4.1) extends to a projective map
+$\mathbb{P}^{1}\longrightarrow\mathbb{P}^{N},\qquad[s:t]\longmapsto[s^{E}:\mathbf{B}(s,t)].$
+We now construct a one-parameter family of lines.
+For simplicity, we assume $N=3$.
+Suppose we have a pair of Bézier curves $\mathbf{B}_{1}(1,t)$ and $\mathbf{B}_{2}(1,t)$ in $\mathbb{K}^{3}$ with degree $E_{1}$ and $E_{2}$, respectively.
+Their projective closures are the rational curves
+$X_{i}:=\left\{[s^{E_{i}}:\mathbf{B}_{i}(s,t)]\in\mathbb{P}^{3}:[s:t]\in\mathbb{P}^{1}\right\}.$
+Under genericity assumptions of the control points, we may assume $X_{1}\cap X_{2}$ is empty. Thus, the points
+$s^{E_{1}}:\mathbf{B}_{1}(s,t)]\text{ and }[s^{E_{2}}:\mathbf{B}_{2}(s,t)]$
+are distinct for all $s:t]\in\mathbb{P}^{1}$, and span a line in $\mathbb{P}^{3}$.
+This set of lines yields a subvariety of $\operatorname{Gr}(1,\mathbb{P}^{3})$, which we denote by $\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}}$.
 This is visualized in Figure 5.
-Again we assume the control points are in generic position.
-The Zariski closure of the family of lines is a subvariety of $\operatorname{Gr}(1,\mathbb{P}^{3})$, and we denote it as $\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}}$.
-The Plücker embedding of this variety is given by the two by two minors of the $2\times 4$ matrix
+Let $\mathcal{A}\subset\mathbb{P}^{11}$ be an irreducible projective variety of $3\times 4$ camera matrices.
+If $C_{1},\dots,C_{n}$ are generic points in $\mathcal{A}$, our line-multiview variety of interest is
+$(C_{1},\dots,C_{n})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}}\;\subset\;\operatorname{Gr}(1,\mathbb{P}^{2})^{n}.$
+Recalling Section 3.1, we say it is anchored at $\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}}$ with respect to the camera arrangement $\mathbf{C}=(C_{1},\dots,C_{n})$.
+Theorem 4.1.
+Let $\mathcal{A}\subset\mathbb{P}^{3\times 4-1}$ be an irreducible projective variety, and fix generic points $C_{1},\dots,C_{n}$ in $\mathcal{A}$.
+For Bézier curves $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$ with $E_{1}+1$ and $E_{2}+1$ generic control points, let $X_{\mathbf{B}_{1},\mathbf{B}_{2},n}$ denote the multiprojective variety
+$\iota\left(\,(C_{1},\dots,C_{n})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}}\,\right)\subset(\mathbb{P}^{2})^{n}.$
+If
+$\operatorname{affEDdeg}(X_{\mathbf{B}_{1},\mathbf{B}_{2},n})=3(E_{1}+E_{2})n-2,$ (4.2)
+holds for $n=1,2$, then (4.2) holds for all $n\geq 1$.
+Proof.
+Under the Plücker embedding a line in $\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}}$ is determined by the six maximal minors of the $2\times 4$ matrix
 $\begin{bmatrix}s^{E_{1}}&\mathbf{B}_{1}(s,t)\\ s^{E_{2}}&\mathbf{B}_{2}(s,t)\end{bmatrix}$.
-This gives a degree $E_{1}+E_{2}$ polynomial map from $\mathbb{P}^{1}$ to $\mathbb{P}^{5}$, and the results of Theorem 2.3 imply
-$\operatorname{affEDdeg}(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}})\,=\,3(E_{1}+E_{2})n-2$
-where $\mathbf{C}$ is an arrangement of $n$ cameras of size $3\times 4$.
-Figure 5: The purple ruled surfaces consist of line segments connecting a pair of Bezier curves of degree one (left) and four (right).
+This yields a degree $E_{1}+E_{2}$ rational map $f_{\mathbf{B}_{1},\mathbf{B}_{2}}:\mathbb{P}^{1}\to\mathbb{P}^{5}$, assuming the genericity of the control points of the Bézier curves.
+By the commutative diagram in Section 3.2, which is a consequence of Section 3.2, we have
+$X_{\mathbf{B}_{1},\mathbf{B}_{2},n}={(D_{1},\dots,D_{n})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\iota(\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}})}$
+where $D_{i}=\wedge^{2}C_{i}$.
+Since ${\iota(\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}})}=f_{\mathbf{B}_{1},\mathbf{B}_{2}}(\mathbb{P}^{1})$ is a rational curve in $\mathbb{P}^{3}$ of degree $E_{1}+E_{2}$, the result follows from Section 2 as the $D_{i}$ are generic points in the irreducible variety $\{\wedge^{2}C:C\in\mathcal{A}\}\subset\mathbb{P}^{3\times 6-1}$.
+Figure 5: The purple ruled surfaces consist of line segments connecting a pair of Bézier curves of degree one (left) and four (right).
 The lines spanned by the segments in a ruling give $\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}}\subset\operatorname{Gr}(1,\mathbb{P}^{3})$.
@@
 We used topological arguments to give the first theoretical results for ED degrees of multiview varieties anchored at a Schubert variety.
-One of the key features of our proof is that each genericity condition is the complement of a union of hypersurfaces whose defining equations only rely on a pair of cameras or a single camera.
+One of the key features of our proof is that each genericity condition is the complement of a union of hypersurfaces whose defining equations only rely on a single camera or a camera pair.
 This resulted in a surprising Section 2 to determine the ED degrees of multiview varieties where cameras have a prescribed structure.
 This is relevant for families of “dual cameras” found in resectioning [6] as well as calibrated cameras [21, 26] used in practice.
-We leave for future work the task of of generalizing these ideas to higher dimensional multiview varieties.
-This becomes more difficult because the camera centers may intersect the world variety.
-We also leave for future work the problem of determining the ED degree of line multiview varieties using a different embedding into the Grassmannian [16, 17, 5].
+A natural extension is to generalize our results on the ED degree of curves to higher-dimensional varieties.
+We leave for future work the task of combining the techniques in [27], which were used to prove (1.5), with the approaches we outlined in our proofs to give a generalization of our Section 2 for all parameterized anchored mutliview varieties.
+Another interesting problem is to determine the ED degree of line multiview varieties using alternative embeddings of the Grassmannian, such as those appearing in recent work on Grassmannian optimization [23, 24, 7].
+A natural next step is to choose a representation of the embedded multiview variety that yields good properties for the bundle adjustment problem.
+For example, we can ask how the study of representations for 3D lines in projective space in [3, Section 3] generalizes to other multiview varieties.
+We hope that further developments along these lines will deepen the interplay between algebraic geometry, multilinear algebra, and computer vision.
+Acknowledgments
+We thank Timothy Duff for his comments on early versions of the article.
+Data availability statement
+The calculations supporting the results of this article can replicated using the open source software Macaulay2 [16] using the EuclideanDistanceDegree.m2 package.
+The code for these calculations is available at the GitHub repository.